### v2

# Junyu Cao and Mohsen Bayati

[…]
Junyu Cao and Mohsen Bayati
[…]
A Probabilistic Approach for Model Alignment with Human Comparisons
[…]
A Probabilistic Approach for Model Alignment with Human Comparisons
[…]
Junyu Cao
[…]
A growing trend involves integrating human knowledge into learning frameworks, leveraging subtle human feedback to refine AI models. While these approaches have shown promising results in practice, the theoretical understanding of when and why such approaches are effective remains limited. This work takes steps toward developing a theoretical framework for analyzing the conditions under which human comparisons can enhance the traditional supervised learning process. Specifically, this paper studies the effective use of noisy-labeled data and human comparison data to address challenges arising from noisy environment and high-dimensional models. We propose a two-stage “Supervised Learning+Learning from Human Feedback” (SL+LHF) framework that connects machine learning with human feedback through a probabilistic bisection approach. The two-stage framework first learns low-dimensional representations from noisy-labeled data via an SL procedure and then uses human comparisons to improve the model alignment. To examine the efficacy of the alignment phase, we introduce a concept, termed the “label-noise-to-comparison-accuracy” (LNCA) ratio. This paper identifies from a theoretical perspective the conditions under which the “SL+LHF” framework outperforms the pure SL approach; we then leverage this LNCA ratio to highlight the advantage of incorporating human evaluators in reducing sample complexity. We validate that the LNCA ratio meets the proposed conditions for its use through a case study conducted via Amazon Mechanical Turk (MTurk).
[…]
model alignment, supervised learning, probabilistic bisection, human-AI interaction.
1 Introduction
In light of the increasing availability of data, the demand for data-driven algorithmic solutions has surged. Machine learning has emerged as the standard technique for numerous tasks, including natural language processing and computer vision, among others. Various streams of operations research also demonstrate its importance.
To model increasingly complex environments, machine learning models have grown in complexity, often leading to over-parameterization. While this increased complexity allows models to capture intricate patterns, it has revealed fundamental limitations in the standard training and testing pipeline. Specifically, the traditional supervised learning (SL) approach of training on noisy-labeled data faces significant challenges due to model misspecification (D’Amour et al. 2022). This issue becomes particularly acute in high-complexity scenarios, such as when the feature dimension is comparable to or exceeds the number of training data points. In such high-dimensional settings, fundamentally different models may achieve similar performance metrics during testing, despite having different underlying structures and behaviors. For instance, as demonstrated by Overman et al. (2024), the trained model that minimizes the empirical risk may violate fundamental principles, such as the natural expectation of negative demand elasticity—indicating that the model is not aligned with its expected behavior. Consequently, decision makers must consider strategic approaches to achieve better model alignment across different environments.
One promising approach has been the incorporation of human feedback into the learning framework (Li et al. 2016, Christiano et al. 2017, Alur et al. 2024). These studies demonstrate that using human feedback for alignment—refining model parameters that are based on human-labeled data—can effectively address the limitations of traditional training approaches. For instance, in the design of chatbots, performance can be adaptively improved through the chatbots’ interactions with humans (Ziegler et al. 2019, Ouyang et al. 2022). Ouyang et al. (2022) exemplify this approach in their work on fine-tuning language models, where InstructGPT is fine-tuned with human feedback. The effectiveness of human feedback stems from evaluators’ domain expertise that extends beyond what can be encoded into a dataset or estimation method. Humans’ deep understanding of the task domain enables them to make informed comparisons that consider context, relevance, and real-world applicability (Deng et al. 2020).
Among various forms of human feedback, comparisons between alternatives have emerged as a particularly effective approach due to their cognitive simplicity. As Carl Jung observes, “Thinking is difficult, therefore let the herd pronounce judgment”—similarly, humans find it easier to select between alternatives than to generate absolute evaluations. Recent work has demonstrated the practical value of such comparative feedback: Zhou and Xu (2020) use pair-wise comparisons for story generation, while Xu et al. (2023) leverage human comparisons between large language models (LLMs) for fine-tuning. Although these applications show promising results, we lack a theoretical framework to understand when and why human comparisons enhance traditional supervised learning. This gap motivates our paper’s central question:
How can we effectively combine supervised learning and human feedback to achieve model alignment?
To make progress towards answering this question, we consider a two-stage probabilistic model that helps us analyze and understand the interplay between supervised learning and human feedback. In the first stage, noisy labeled data are fed into the SL oracle to obtain low-dimensional representations (or embeddings). For example, selecting a small set of features through Lasso is a way to represent high-dimensional feature spaces; similarly, low-rank matrix estimation is an example of representation of high-dimensional matrices. The second stage uses human comparisons to improve model alignment within this low-dimensional space. Our framework uses the random utility model to describe humans’ comparison behavior. Practically, humans may provide incorrect responses, so we develop a probabilistic bisection approach for the alignment phase. We call this two-stage process the Supervised Learning+Learning from Human Feedback (SL+LHF) model.
The rationale behind this model is rooted in compelling evidence that representation learning techniques, such as deep learning architectures (Zhou and Paffenroth 2017, Akrami et al. 2019), can effectively learn low-dimensional features even from highly noisy data (Li et al. 2021, Taghanaki et al. 2021, Mousavi-Hosseini et al. 2024). For instance, Tsou et al. (2023) recently showed that linear probes trained on the final embedding of pre-trained GPT-2 are surprisingly robust, sometimes even outperforming full-model fine-tuning. However, while supervised learning excels at learning these representations, the challenge lies in effectively fine-tuning the final layer—the stage where the model’s predictions are aligned with desired behaviors. For this alignment phase, we leverage human expertise through comparisons rather than direct labels, building on the cognitive advantage of comparative judgments discussed above. Our approach uses a sequential design of comparisons and active data acquisition to maximize information gain and minimize sample complexity, particularly effective when label noise is high.
To describe the effectiveness of the refinement phase involving human comparison, we introduce a new concept called the “label-noise-to-comparison-accuracy” (LNCA) ratio. We offer a theoretical characterization of the condition in which the SL+LHF framework performs better than the pure SL algorithm based on this LNCA ratio. In a large-noise scenario, human evaluators with high selection accuracy can significantly reduce the sample complexity.
1.1 Our Contributions
This paper makes contributions in four areas: theoretical modeling, characterization of sample complexity, practical implementation, and empirical validation.
Modeling.
We develop a theoretical framework that explains when and why human comparisons can substantially enhance performance of SL in high-noise environments. Our model captures the interaction between SL and human feedback through a utility-based comparative judgment approach.
Theoretical contributions.
Our main theoretical results consist of two components. First, we develop a probabilistic bisection method—an efficient sequential decision-making process that alternates between “vertical moves” (multiple evaluations of the same comparison) and “horizontal moves” (selection of new comparison pairs) to acquire human comparisons. Second, focusing on settings where the SL stage utilizes Lasso for learning low-dimensional representations, we introduce the label noise-to-comparison-accuracy ratio (LNCA) that quantifies the trade-off between label noise and human comparison accuracy. Our goal is to refine the model within an $\varepsilon$-distance with confidence at least $1-\delta$, which $\varepsilon$ represents the learning precision and $1-\delta$ denotes the confidence level. When human selection accuracy uniformly exceeds random guessing, we prove the refinement stage has sample complexity $O(s\log(1/(\delta\varepsilon)))$, where $s$ is the number of non-zero coefficients in Lasso. We characterize conditions under which our SL+LHF framework outperforms pure SL.
Towards practical implementations.
While our theoretical analysis examines human comparisons of model coefficients for analytical tractability, direct coefficient comparison is impractical in real applications. To bridge this gap, we develop an Active Sample Selection (ASS) framework that translates our theoretical insights into implementable procedures by enabling humans to compare meaningful data points rather than abstract model parameters. This framework provides a systematic approach for selecting the most informative samples for human comparison, applicable across a broad set of data distributions.
Empirical analysis.
Through Amazon Mechanical Turk (MTurk) experiments to acquire real human feedback on a stylized task, we validate our theoretical findings about the LNCA ratio and demonstrate the superior performance of SL+LHF over pure SL under conditions of higher observational noise, higher human comparison accuracy, and lower-dimensional feature spaces.
The remainder of this paper is organized as follows: Section 2 reviews related literature. Section 3 presents our model setup, followed by Section 4, which introduces human comparison modeling and characterizes sample complexity in one dimension. Section 5 develops our two-stage framework integrating SL and human comparisons, with particular focus on sample complexity when using Lasso in the first stage. Section 6 details practical implementations for active sample selection, while Section 7 presents numerical experiments and managerial insights. Section 8 concludes. All proofs and parameter calibrations appear in the appendices.
2 Literature Review
Alignment and Learning from Human Feedback. The alignment challenge seeks to synchronize human values with machine learning systems and direct learning algorithms toward the objectives and interests of humans. Machine learning models frequently demonstrate unforeseen deficiencies in their performance when they are implemented in practical contexts; this phenomenon has been identified as underspecification (D’Amour et al. 2022). Predictors that perform similarly in the training dataset can exhibit significant variation when they are deployed in other domains. The presence of ambiguity in a model might result in instability and suboptimal performance in real-world scenarios. Recent papers have focused on aligning models with human intentions, including training simple robots in simulated environments and Atari games (Christiano et al. 2017, Ibarz et al. 2018) and fine-tuning language models to summarize text (Stiennon et al. 2020) and to optimize dialogue (Jaques et al. 2019). Studies have demonstrated that using reinforcement learning from human feedback (RLHF) significantly boosts performance (Bai et al. 2022, Glaese et al. 2022, Stiennon et al. 2020, Dwaracherla et al. 2024). Ouyang et al. (2022) use RLHF (Stiennon et al. 2020) to fine-tune GPT-3 to follow a broad class of written instructions. In particular, Zhou and Xu (2020) use pair-wise comparisons for story generation. They find that, when comparing two natural language generation (NLG) models, asking a human annotator to assign scores separately for samples generated by different models, which resembles the case in the ADEM model (Lowe et al. 2017), is more complicated. The much easier approach is for human annotators to directly compare one sample generated by the first model against another sample from the second model in a pair-wise fashion and then to compute the win/loss rate. Zhu et al. (2023) provide a reinforcement learning framework that includes human feedback – specifically, for function approximation in $K$-wise comparisons, with policy learning as the target. Xu et al. (2023) propose a fine-tuning algorithm by inducing a complementary distribution over the two sampled LLM responses to model human comparison processes, in accordance with the Bradley-Terry model for pairwise comparisons. In contrast to all these works, we provide a theoretical, two-stage framework for integrating supervised learning with actively acquired human comparison labels, and we characterize the conditions where the two-stage framework outperforms pure SL algorithms.
Transfer Learning and Task Alignment. Model alignment generally refers to the process of ensuring that machine learning models are tailored to meet specific tasks and objectives. A related paradigm is transfer learning, where a model trained on one task is adapted for use in another task through additional fine-tuning (Pan and Yang 2009, Xu et al. 2021, Bastani et al. 2022, Du et al. 2024). While our work shares some conceptual similarities with transfer learning, we focus specifically on the impact of noise in both supervised learning and human feedback phases. Our framework’s applicability to transfer learning scenarios depends crucially on the relationship between the initial and target tasks’ representations. When the target task shares the same low-dimensional representation as the initial task, requiring only updates to the final layer or coefficients, our theoretical results apply directly—the human comparison phase can effectively refine these parameters regardless of whether the comparison objective exactly matches the initial task. However, when the target task requires different representations, transfer learning typically demands an additional supervised learning phase with new labeled data to first refine these representations before any human-guided fine-tuning can begin. In such cases, our framework’s human comparison stage remains relevant for the final alignment when label noise is high, but it must be preceded by this representation learning phase.
Active Learning. The core concept behind active learning in machine learning theory is the ability of a model to interactively query a human or an oracle to obtain labels for new data points. Instead of passively learning from a fixed dataset, an active learning algorithm intelligently selects the most informative learning instances to query for labels. We refer readers to Settles (2009) for a comprehensive review of the active learning algorithm literature. The literature has concentrated on studying label complexity to demonstrate the benefit of active learning, which refers to the number of labels requested to attain a specific accuracy. For example, Cohn et al. (1994) demonstrate that active learning can have a substantially lower label complexity than supervised learning in the noiseless binary classification issue. In the presence of noise, other papers focus on the classification problems (Hanneke 2007, Balcan et al. 2006, Hanneke 2011) and the regression problems (Sugiyama and Nakajima 2009, Beygelzimer et al. 2009). Although our study has a similar goal of using the fewest possible samples to achieve the target accuracy, we concentrate on the human-AI collaboration framework and try to reduce the sample complexity of human comparisons.
Data-Driven Decision Making and Humans in the Loop. A large body of recent literature on operations research or operations management has studied how to effectively use data to improve decisions (Mišić and Perakis 2020), including pricing, assortment, and recommendation. Motivated by the observation that humans may sometimes have “private” information to which an algorithm does not have access but that can improve performance, Balakrishnan et al. (2022) examine new theory that describes algorithm overriding behavior. They designed an experiment to test feature transparency as an implementable approach that can help people better identify and use their private information. Ibrahim et al. (2021) study how to address the shared information problem between humans and algorithms in a setting where the algorithm makes the final decision using the human’s prediction as input. Dietvorst et al. (2018, 2015) study the algorithm aversion phenomenon, in which people fail to use algorithms after learning that they are imperfect, even though evidence-based algorithms consistently outperform human forecasters. Other papers study human-artificial intelligence (AI) collaboration on a large variety of topics (Bastani et al. 2021, Wu et al. 2022, Deng et al. 2020, Chen et al. 2022, Caro and de Tejada Cuenca 2023, Ye et al. 2024), including ride-hailing (Benjaafar et al. 2022) and health care (Reverberi et al. 2022).
Probabilistic Bisection. Another stream of literature related to our work, which is independent of the previously identified human-AI ML topics, is on probabilistic bisection algorithms (PBAs) (Horstein 1963). Algorithms using the PBA framework have been used in several contexts – notably, including the tasks of target localization (Tsiligkaridis et al. 2014), scanning electron microscopy (Sznitman et al. 2013), topology estimation (Castro and Nowak 2008), edge detection (Golubev and Levit 2003), and value function approximation for optimal stopping problems (Rodriguez and Ludkovski 2015). From a computational perspective, Frazier et al. (2019) have shown that probabilistic bisection converges almost as quickly as stochastic approximation. In this work, we design the model alignment algorithm by using the PBA approach, where randomness is involved because of the inherent variability in the human judgment process.
3 Model
In this section, we introduce our model setup. We first introduce the supervised learning oracle that potentially is used during the initial learning stage, and then we present the fine-tuning procedure involving human feedback during the refinement stage. The two alternative ways of collecting human feedback are either to collect human labels directly or to ask humans to compare two answers. The former approach, which continues the label collecting procedure following the initial learning stage, is called pure supervised learning. As previously noted, we call the latter approach the Supervised Learning+Learning from Human Feedback framework.
3.1 Supervised Learning and Underspecifications
Given the data $\{(X_{i},Y_{i})\}_{i\leq n}$, where $X_{i}\in\mathcal{X}$ and $Y_{i}\in\mathcal{Y}\subseteq\mathbb{R}$, the goal is to learn the true function that is parameterized by $\boldsymbol{\vartheta}^{*}\subseteq\mathbb{R}^{d}$, $h_{\boldsymbol{\vartheta}^{*}}\in\mathcal{H}:\mathcal{X}\rightarrow\mathcal{Y}$, where $\mathcal{H}$ is the function class. Each data point is generated from a stochastic response:
$Y_{i}=h_{\boldsymbol{\vartheta}^{*}}(X_{i})+\epsilon_{i}\,,$
where $\epsilon_{i}$ is the noise with mean 0 and variance $\sigma^{2}$, and where $\sigma$ could potentially be very large.
We specifically consider a high-dimensional problem that can be represented by the low-dimensional generalized linear model; that is,
$h_{\boldsymbol{\vartheta}^{*}}(x)=f(\left\langle\boldsymbol{\theta}^{*},\varphi(x)\right\rangle)\,,$
where $f(\cdot)$ is a link function. Here, $\boldsymbol{\theta}^{*}\in\Theta\in\mathbb{R}^{s}$ is the unknown parameter [^1] We denote the parameter vector using the bold symbol.; and $\varphi(x)\in\mathbb{R}^{s}$ is the kernel function, which is the low-dimensional representation of feature $x$. Our goal is to learn both the parameter $\boldsymbol{\theta}^{*}$ and representation $\varphi(x)$. We assume that $s\ll d$.
We provide the following two examples as illustrations.
Example 3.1 (Sparse Linear Models)
Suppose $x$, $\boldsymbol{\vartheta}^{*}\in\mathbb{R}^{d}$, and $h_{\boldsymbol{\vartheta}^{*}}(x)=\left\langle\boldsymbol{\vartheta}^{*},x\right\rangle$. The parameter $\boldsymbol{\vartheta}^{*}$ has nonzero entries in the first $s$ dimensions and has zero entries in the rest of the dimensions. In this case, $h_{\boldsymbol{\vartheta}^{*}}(x)=f(\left\langle\boldsymbol{\theta}^{*},\varphi(x)\right\rangle)$ where $\boldsymbol{\theta}^{*}=(\vartheta^{*}_{1},\cdots,\vartheta^{*}_{s})$, $\varphi(x)=(x_{1},\cdots,x_{s})$, and $f(\cdot)$ is an identity function.
Example 3.2 (Generalized Low-rank Models)
Suppose $x\in\mathbb{R}^{d\times d}$. In the generalized low-rank models, we assume that $h_{\boldsymbol{\vartheta}^{*}}(x)=f(\left\langle\boldsymbol{\Theta}^{*},x\right\rangle)$ where $\boldsymbol{\Theta}^{*}\in\mathbb{R}^{d\times d}$ is a low-rank matrix (i.e., $\boldsymbol{\Theta}^{*}=\sum_{i=1}^{s}\sigma_{i}u_{i}v_{i}^{\top}$ where $u_{i}$ and $v_{i}\in\mathbb{R}^{d})$; the trace inner product is defined as $\left\langle\boldsymbol{\Theta}^{*},x\right\rangle=tr(\boldsymbol{\Theta}^{*}x^{\top})$. Thus,
$h_{\boldsymbol{\vartheta}^{*}}(x)=f(\sum_{i=1}^{s}\sigma_{i}\cdot tr(u_{i}v_{i}^{\top}x^{\top}))=f(\left\langle\boldsymbol{\theta}^{*},\varphi(x)\right\rangle),$
where $\varphi(X)=[tr(u_{i}v_{i}^{\top}x^{\top})]_{i=1}^{s}$ is the projection of $x$ onto the low-rank space, and $\boldsymbol{\theta}^{*}=(\sigma_{1},\cdots,\sigma_{s})$ is the $s$-dimensional vector of singular values.
Suppose that $h_{\boldsymbol{\vartheta}}$ can be estimated through a statistical loss function $\ell:\mathcal{Y}\times\mathbb{R}\rightarrow\mathbb{R}$ and that the total loss is
$L_{n}(h_{\boldsymbol{\vartheta}})=\sum_{i=1}^{n}\ell(Y_{i},h_{\boldsymbol{\vartheta}}(X_{i})).$
For example, $\ell$ can represent the squared loss and the regularized squared loss, among others. To deal with the high-dimensional issue, the algorithm learns the low-dimensional representation $\varphi(x)$, as well as the parameter $\boldsymbol{\theta}^{*}$. For example, to learn the sparse linear models, one can apply Lasso to select important features and to learn the model parameters. To learn the low-rank models, one can use trace regression to learn the low-dimensional representations.
Define $h_{\hat{\boldsymbol{\vartheta}}_{n}}$ as the function contained in $\mathcal{H}$ that minimizes empirical risk $L_{n}$. In the existing literature, the decision maker very commonly employs the predictor $h_{\hat{\boldsymbol{\vartheta}}_{n}}$ directly. However, some issues may arise. In high-dimensional problems, where significant noise is present, approaches striving to balance the variance–bias trade-off may violate fundamental principles, such as the natural expectation of negative demand elasticity in economics. We use the following example to illustrate this potential issue.
Example 3.3
Consider a linear pricing model with high-dimensional contexts – that is, $Y_{i}=\vartheta_{1}\cdot p_{i}+\vartheta_{2}^{\top}X_{i}+\epsilon_{i}$, where $Y_{i}$ is the demand, $p_{i}$ is the price, and the vector $X_{i}$ represents features of product $i$. Also, assume that $\vartheta_{1}<0$. However, in a high-noise regime, we may incorrectly estimate a positive value for $\vartheta_{1}$. Please see Appendix 10 for numerical illustrations.
In Example 3.3, the negative coefficient of the price in the demand function is misinterpreted because of model underspecification. The requirement of $\vartheta_{1}<0$ should be specified based on human knowledge. In the face of the challenge that underspecification poses in a potentially very noisy environment, we take a critical step toward refinement; we call this step “model alignment through human comparisons.” Humans can do comparisons with much greater precision. For example, on the basis of their knowledge, humans can immediately correct the negative correlation between demand and price. This particular stage comes into play after our machine learning pipelines generate and produce their output $\hat{\varphi}(\cdot)$ and $\hat{\boldsymbol{\theta}}_{0}$ (which denotes the estimator of $\varphi$ and $\boldsymbol{\theta}$) by training on the noisy-labeled data. During the second stage, we refine the model in a lower dimensional space. Throughout the paper, we focus our analysis on sparse linear models.
3.2 Utility Model of Human Comparisons
In this section, we introduce the human comparison model. Given two potential models, $f_{\boldsymbol{\theta}_{1}}$ and $f_{\boldsymbol{\theta}_{2}}$, the system asks humans to compare and select the better prediction model; we call this step the model-level comparison. If $f_{\boldsymbol{\theta}_{1}}$ is preferable to $f_{\boldsymbol{\theta}_{2}}$, we denote the result as $f_{\boldsymbol{\theta}_{1}}\succeq f_{\boldsymbol{\theta}_{2}}$.
Remark 3.4 (Model-level Comparison and Sample-level Comparison)
In practice, the comparison between two models can be conducted in various ways. For example, the models can be evaluated on a single data point (either from the humans’ past experience or their observation), denoted as $\{x,y\}$. When comparing two models in this case, the one with a smaller squared loss is preferred; that is:
$\mathbb{P}(\boldsymbol{\theta}_{1}\succeq\boldsymbol{\theta}_{2})$ $=\mathbb{P}\left((y-f_{\boldsymbol{\theta}_{1}}(x))^{2}\leq(y-f_{\boldsymbol{\theta}_{2}}(x))^{2}\right).$
Alternatively, models also can be assessed on a batch of samples. We call this approach the sample-level comparison. Note that a model-level comparison is a broader concept. We elaborate on how to translate the model-level comparison to the sample-level comparison in Section 6. In this section, our focus is on a framework for undertaking a more general model-level comparison.
The preference over the model $f_{\boldsymbol{\theta}}$ is captured by a utility function $u(\boldsymbol{\theta})$, where $u(\boldsymbol{\theta})$ depicts the distance between the model parameterized by $\boldsymbol{\theta}$ and the true model. That is, we assume $u(\boldsymbol{\theta})=-d(\boldsymbol{\theta},\boldsymbol{\theta}^{*})$, where $d(\cdot,\cdot)$ is a distance function. The larger utility value indicates the better model. The true model parameter $\boldsymbol{\theta}^{*}$ achieves the highest utility. For example, it can be the negative two-norm distance – that is, $u(\boldsymbol{\theta})=-\|\boldsymbol{\theta}-\boldsymbol{\theta}^{*}\|_{2}$. We formally state the assumption about the utility function as follows:
{assumption}
[Utility function] Assume $u(\boldsymbol{\theta})=-d(\boldsymbol{\theta},\boldsymbol{\theta}^{*})$, where $d(\cdot,\cdot)$ is a distance function. The utility function is unimodal in each coordinate. That is, for each coordinate $\theta_{i}$, when fixing all other coordinates, $u(\boldsymbol{\theta})$ is monotonically increasing when $\theta_{i}<\theta_{i}^{*}$ and is monotonically decreasing when $\theta_{i}>\theta_{i}^{*}$. Moreover, the utility gets lower when $\boldsymbol{\theta}$ gets far away from the true parameter. The utility function can be either bounded or unbounded (e.g., the Euclidean distance). We give another example of the bounded utility function as an illustration.
Example 3.5
[…]
The comparison between $f_{\boldsymbol{\theta}_{1}}$ and $f_{\boldsymbol{\theta}_{2}}$ provides information on which side is closer to the true parameter $\boldsymbol{\theta}^{*}$. Define the hyperplane $H(\boldsymbol{\theta}_{1},\boldsymbol{\theta}_{2})=\{\boldsymbol{\theta}:d(\boldsymbol{\theta},\boldsymbol{\theta}_{1})=d(\boldsymbol{\theta},\boldsymbol{\theta}_{2})\}$, as well as half-space $H^{-}(\boldsymbol{\theta}_{1},\boldsymbol{\theta}_{2})=\{\boldsymbol{\theta}:d(\boldsymbol{\theta},\boldsymbol{\theta}_{1})<d(\boldsymbol{\theta},\boldsymbol{\theta}_{2})\}$ and half-space $H^{+}(\boldsymbol{\theta}_{1},\boldsymbol{\theta}_{2})=\{\boldsymbol{\theta}:d(\boldsymbol{\theta},\boldsymbol{\theta}_{1})>d(\boldsymbol{\theta},\boldsymbol{\theta}_{2})\}$. The condition $u(\boldsymbol{\theta}_{1})\geq u(\boldsymbol{\theta}_{2})$ implies that $\boldsymbol{\theta}^{*}\in H^{-}(\boldsymbol{\theta}_{1},\boldsymbol{\theta}_{2})$. Assumption 3.4 implies that $H(\theta_{1},\theta_{2})$ contains a unique point when $\theta_{1}$ and $\theta_{2}$ are in a one-dimensional space.
The overall goal is to find the $\boldsymbol{\theta}$ that achieves near the highest utility or that, equivalently, gets close to the true model parameter $\boldsymbol{\theta}^{*}$ within a certain precision. Suppose that the human selection process follows a random utility model (RUM). In the process of querying, the human’s opinion regarding $f_{\boldsymbol{\theta}}$ is
$U(\boldsymbol{\theta})=u(\boldsymbol{\theta})+\xi,$ (3.1)
where $\xi$ is the unobserved noise. We assume that the noise $\xi$ follows a Gumbel distribution with parameter $\gamma$, which represents the expert level of the human. We say human evaluators make the correct choice if they select $\boldsymbol{\theta}_{1}$ ($\boldsymbol{\theta}_{2}$) when $u(\boldsymbol{\theta}_{1})$ ($u(\boldsymbol{\theta}_{2})$) is higher than $u(\boldsymbol{\theta}_{2})$ ($u(\boldsymbol{\theta}_{1})$). A smaller value of $\gamma$ corresponds to a higher certainty of the selection’s correctness. When $\gamma$ approaches 0, humans can select the better one with almost 100% accuracy.
Remark 3.6
Our model can also incorporate the human bias in the following way: $U(\boldsymbol{\theta})=u(\boldsymbol{\theta})+\beta_{h}+\xi$, where the human bias $\beta_{h}$ can vary between individuals. For simplicity, we focus on the utility model (3.1).
Lemma 3.7 (Precision)
For any two models $f_{\boldsymbol{\theta}_{1}}$ and $f_{\boldsymbol{\theta}_{2}}$, the probability that a human will make the right selection is
$\frac{1}{1+\exp(-|u(\boldsymbol{\theta}_{2})-u(\boldsymbol{\theta}_{1})|/\gamma)}\,,$
which is strictly greater than 1/2.
Lemma 3.7 implies that under a RUM, the selection accuracy is always higher than 50%, even with the nonzero human bias. Intuitively, the differentiation would become harder when $\boldsymbol{\theta}_{1}$ and $\boldsymbol{\theta}_{2}$ both get closer to the true parameter. However, how fast it converges to 1/2 when $\boldsymbol{\theta}_{1}$ and $\boldsymbol{\theta}_{2}$ approach $\boldsymbol{\theta}^{*}$ depends on the distance function. Under this probabilistic selection model, we design the algorithm and investigate the sample complexity. The parameter $\boldsymbol{\theta}$ may lie in a multi-dimensional space. To illustrate our algorithm, we first describe the one-dimensional human comparison procedure, and then we extend to the multi-dimensional scenario.
4 One-dimensional Human Comparison
The low-dimensional representation $\hat{\varphi}$ is established in the initial stage, as we discuss in Section 5. Following this stage, our framework calls for model alignment through human comparisons. At this point, the parameter $\boldsymbol{\theta}$ may lie in a multi-dimensional space. To illustrate our algorithm, we first describe the one-dimensional human comparison procedure and then extend it to the multi-dimensional scenario.
The primary challenge in designing the framework lies in strategically selecting pairwise comparisons to minimize the number of comparisons (or sample complexity), which resembles the principles of active learning. First, we consider a special case in which humans can select the better model with 100% accuracy (deterministic selection). Second, we relax this to a probabilistic selection. For the probabilistic selection, we propose an algorithm we call Model Alignment through Probabilistic Bisection (MAPB) and characterize its sample complexity.
4.1 Deterministic Bisection
We first illustrate the algorithm in a one-dimensional space with 100% selection accuracy. This scenario corresponds to $\gamma\to 0$ in RUM. In the one-dimensional case, suppose $\theta^{*}\in\Theta\subseteq\mathbb{R}$ and $|\theta|\leq\beta_{\Theta}$ for all $\theta\in\Theta$.
The overall goal is to use the least number of samples to find a $\theta$ such that $d(\theta,\theta^{*})\leq\varepsilon$. When a human is able to select a better model with 100% accuracy, the deterministic bisection algorithm can be used to approach the true parameter. The algorithm operates by maintaining a search interval $[\theta^{-},\theta^{+}]$ that contains the true parameter $\theta^{*}$. The interval is initialized by $[-\beta_{\Theta},\beta_{\Theta}]$, and at round $k$, the center of this interval with respect to the distance $d(\cdot,\cdot)$ is used as the query point. Specifically, after defining the query point $\theta_{k}$ to be the single point in the set $H(\theta^{-},\theta^{+})$, we ask the human to compare the utility of two points, denoted by $c_{\Delta}^{-}(\theta_{k})$ and $c_{\Delta}^{+}(\theta_{k})$. These two points have equal distance $\Delta/2$ from $\theta_{k}$ in opposite directions, where $\Delta$ is called comparison granularity and is a tuning parameter. For example, for the Euclidean distance, at each round, $\theta_{k}=(\theta^{-}+\theta^{+})/2$; also, $c_{\Delta}^{-}(\theta_{k})=\theta_{k}-\Delta/2$, and $c_{\Delta}^{+}(\theta_{k})=\theta_{k}+\Delta/2$. If the human selects $c_{\Delta}^{-}(\theta_{k})$, then we eliminate the upper half of the search interval by updating $\theta^{+}$ to be equal to $\theta_{k}$; otherwise, we eliminate the bottom half of the search interval by setting $\theta^{-}$ to be equal to $\theta_{k}$. After this bisection step, we continue the process by selecting $\theta_{k+1}$ to be $H(\theta^{-},\theta^{+})$ again, and repeat the above step until the length of the search interval falls below the precision $\varepsilon$. The pseudo-code is presented in Algorithm 1.
$\theta_{k}$$\theta^{*}$$\theta^{-}$$\theta^{+}$$c_{\Delta}^{+}(\theta_{k})$$c_{\Delta}^{-}(\theta_{k})$keepeliminate
Figure 1: An illustrative example. Among two choices $c_{\Delta}^{-}(\theta_{k})$ and $c_{\Delta}^{+}(\theta_{k})$, $c_{\Delta}^{-}(\theta_{k})$ would be selected because it is closer to the true parameter. In this case, the interval $(\theta_{k},\theta^{+}]$ is eliminated.
Algorithm 1 DB (Deterministic Bisection)
1:input: precision $\varepsilon$; comparison granularity $\Delta$; $k=0$; $\theta_{k}=0$; $\theta^{+}=\beta_{\Theta}$; $\theta^{-}=-\beta_{\Theta}$;
2:repeat
3:     Ask human to evaluate $c_{\Delta}^{-}(\theta_{k})$ and $c_{\Delta}^{+}(\theta_{k})$
4:     if $c_{\Delta}^{+}(\theta_{k})$ is chosen then
5:         $\theta^{-}=\theta_{k}$;
6:     else $\theta^{+}=\theta_{k}$;
7:     end if
8:     $k=k+1$; $\theta_{k}=H(\theta^{-},\theta^{+})$;
9:until $d(\theta^{+},\theta^{-})\leq\varepsilon$.
We can show the following result for the performance of Algorithm DB.
Proposition 4.1
Suppose the human makes the selection with 100% accuracy $(\gamma\to 0)$. Fix $\varepsilon>0$. After $O(\log_{2}\left(\frac{\beta_{\Theta}}{\varepsilon}\right))$ rounds, Algorithm 1 reaches a point $\theta_{k}$ such that $d(\theta_{k},\theta^{*})\leq\varepsilon$.
Note that the comparison granularity has no impact on Algorithm 1 and its computational complexity because the deterministic model assumes 100% accuracy for any $\Delta$. However, this deterministic model of human selection as described exhibits two significant limitations: First, in practical scenarios, humans may not always be able to identify the correct choice, implying that $\gamma>0$; second, the accuracy of human selections is influenced by the comparison granularity $\Delta$. To address these limitations, we introduce a more general probabilistic model in the following section.
4.2 Probabilistic Bisection
First, we enrich the model to capture the probabilistic nature of human comparisons by noting that, in the RUM, when $\gamma$ is strictly positive, the human’s selection is random, with a certain distribution that favors the correct answer, parameterized by $\gamma$. Precisely, as shown in Lemma 3.7, the human selects $c_{\Delta}^{-}(\theta)$ with probability
$p^{-}_{\gamma}(\theta):=\left[1+\exp\left\{\frac{u\Big{(}c_{\Delta}^{+}(\theta)\Big{)}-u\Big{(}c_{\Delta}^{-}(\theta)\Big{)}}{\gamma}\right\}\right]^{-1}\,,$
and the human selects $c_{\Delta}^{+}(\theta)$ with probability $p^{+}_{\gamma}(\theta)$, which is equal to $1-p^{-}_{\gamma}(\theta)$.
We note that, in this equation, the selection probability depends on the comparison granularity. This dependence naturally captures the difficulty of the selection for the human. On the one hand, when $\Delta$ is too small, $c_{\Delta}^{-}(\theta)$ and $c_{\Delta}^{+}(\theta)$ converge toward each other; hence, $p^{-}_{\gamma}(\theta)$ and $p^{+}_{\gamma}(\theta)$ converge to $1/2$, representing the difficulty of distinguishing between two highly similar choices. On the other hand, when $\Delta$ is too large, the selection can become difficult as well because both options are far away from $\theta^{*}$. To illustrate, this difficulty can be captured when the distance function converges to a finite value for far away points from $\theta^{*}$, as in Example 3.5. In this case, $p^{-}_{\gamma}(\theta)$ and $p^{+}_{\gamma}(\theta)$ could also converge to 1/2 (see Example 4.2). As a result of this phenomenon, and for brevity, we exclude $\Delta$ in the notations $p^{-}_{\gamma}(\theta)$ and $p^{+}_{\gamma}(\theta)$ in the remainder of the paper.
Example 4.2
Consider
$u(\theta)=-\frac{1}{1+\exp(-|\theta|)}\,.$
Here, the highest utility is achieved at $\theta=0$. For any finite $\theta$, when $\Delta$ is too large, both $u(c_{\Delta}^{+}(\theta))$ and $u(c_{\Delta}^{-}(\theta))$ are close to -1, and thus $p^{-}_{\gamma}(\theta)$ and $p^{+}_{\gamma}(\theta)$ are both close to 1/2.
Definition 4.3 ($(\varepsilon,\delta)$-alignment problem)
Let $\varepsilon>0$ denote the precision level and let $\delta>0$ represent the confidence level. The $(\varepsilon,\delta)$-alignment problem aims to identify a $\theta$ such that $\mathbb{P}(d(\theta,\theta^{*})\leq\varepsilon)\geq 1-\delta$.
Our objective is to solve the $(\varepsilon,\delta)$-alignment problem with small sample complexity. By selecting appropriate values for these parameters, we can also control other metrics such as risk. The algorithm design is inspired by the probabilistic bisection developed by Frazier et al. (2019). Although their work primarily focuses on the analysis of convergence rates, our emphasis is on conducting a sample complexity analysis for an algorithm designed to terminate when both the precision level and confidence level criteria are met.
If $\theta<\theta^{*}$, then $p_{\gamma}^{-}(\theta)<1/2$ and $p_{\gamma}^{+}(\theta)>1/2$; and vice versa. We introduce several notations to enhance simplicity. Let $Y(\theta)$ denote the selection: $Y(\theta)=c_{\Delta}^{-}(\theta)$ with probability $p_{\gamma}^{-}(\theta)$ and $Y(\theta)=c_{\Delta}^{+}(\theta)$ with probability $p_{\gamma}^{+}(\theta)$. Define $\tilde{Z}(\theta)=2\cdot\boldsymbol{1}(Y(\theta)=c_{\Delta}^{+}(\theta))-1$, so that $\tilde{Z}(\theta)=1$ if $Y(\theta)=c_{\Delta}^{+}(\theta)$ (when we believe $\theta^{*}$ is more likely to be to the right of $\theta$) and $\tilde{Z}(\theta)=-1$ if $Y(\theta)=c_{\Delta}^{-}(\theta)$ (when we believe $\theta^{*}$ is more likely to be to the left of $\theta$). Note that we have $\tilde{Z}(\theta)=1$ with probability $p_{\gamma}^{+}(\theta)$. Also, define
$\tilde{p}(\theta)=\mathbb{P}\big{(}\tilde{Z}(\theta)\text{ correctly identifies the direction}\big{)}=\begin{cases}\mathbb{P}(\tilde{Z}(\theta)=1),&\mbox{if }\theta\leq\theta^{*}\\ \mathbb{P}(\tilde{Z}(\theta)=-1),&\mbox{if }\theta>\theta^{*}\end{cases}\,.$
Figure 2 shows the structure of introducing our main algorithm. Algorithm MAPB, described as follows, involves two types of moves: the horizontal moves and the vertical moves. For any fixed $\theta$, we ask the human to repeatedly evaluate $c_{\Delta}^{-}(\theta)$ and $c_{\Delta}^{+}(\theta)$. We refer to this as the vertical moves. After the vertical stopping criteria are satisfied, the algorithm decides the next $\theta$ to evaluate according to a certain distribution that will be defined. We call this step the horizontal move. The horizontal moves continue until the horizontal stopping criteria are satisfied. In the following two sections, we discuss the vertical moves and horizontal moves, as well as their stopping criteria, in sequence. Then, as shown in Figure 2, we introduce the two-stage Human-AI collaboration framework for multiple dimensions in Section 5 and practical implementations in Section 6.
One-dimensional MAPB (Section 4.2) Horizontal moves (Section 4.2.2); Complexity characterization (Section 4.3) Sample complexity (Theorem 4.5) Extension to multi-dimensional setting (Section 5); Two-stage framework for human-AI collaboration (Section 5.1); Complexity characterization (Section 5.3) Practical implementations (Section 6) Vertical moves (Section 4.2.1); Complexity characterization (Section 4.4)
Figure 2: Roadmap for introducing Algorithm 2.
4.2.1 Vertical moves
Fix some $\theta$. We repeatedly ask for human evaluations of the query point $\theta$, achieved through comparisons of $c^{+}_{\Delta}(\theta)$ and $c^{-}_{\Delta}(\theta)$. We observe a sequence of signs $\{\tilde{Z}_{i}(\theta)\}_{i}$. If $\theta<\theta^{*}$, then the expectation $\mathbb{E}[\tilde{Z}_{i}(\theta)]=2\tilde{p}(\theta)-1>0$; if $\theta>\theta^{*}$, then the expectation $\mathbb{E}[\tilde{Z}_{i}(\theta)]=1-2\tilde{p}(\theta)<0$. From the response $\tilde{Z}_{i}(\theta)$, we define the simple random walk $\{S_{m}(\theta):m\geq 0\}$ with $S_{0}(\theta)=0$ and $S_{m}(\theta)=\sum_{i=1}^{m}\tilde{Z}_{i}(\theta)$ for $m\geq 1$. Sequential test of power one indicates whether the drift $\omega$ of a random walk satisfies the hypothesis $\omega<0$ versus $\omega>0$.
Define the stopping time $\tau^{\uparrow}_{1}(\theta)=\inf\{m\in\mathbb{N}:|S_{m}(\theta)|\geq\hbar_{m}\}$, where $\{\hbar_{m}\}_{m}$ is a positive sequence. Here, $\uparrow$” and $\rightarrow$” represent vertical and horizontal, respectively. The test decides that $\theta<\theta^{*}$, if $S_{\tau_{1}^{\uparrow}(\theta)}(\theta)\geq\hbar_{\tau^{\uparrow}_{1}(\theta)}$; it decides that $\theta>\theta^{*}$ if $S_{\tau_{1}^{\uparrow}(\theta)}(\theta)\leq-\hbar_{\tau^{\uparrow}_{1}(\theta)}$; and the test does not make a decision if $\tau^{\uparrow}_{1}(\theta)=\infty$.
Assume now that the algorithm queries a point $\theta_{k}$ (not equal to $\theta^{*}$) at the $(k+1)^{th}$ iteration. We then observe a random walk with the $m^{th}$ term $S_{k,m}=S_{k,m}(\theta_{k})=\sum_{i=1}^{m}Z_{k,i}(\theta_{k})$ until we reach the end of the test, which is called the power-one test (Robbins and Siegmund 1974). Define the new signal:
$Z_{k}(\theta_{k})=\begin{cases}+1,&\mbox{if }S_{k,\tau_{1}^{\uparrow}(\theta_{k})}>0\,,\\ -1,&\mbox{if }S_{k,\tau_{1}^{\uparrow}(\theta_{k})}<0\,.\end{cases}$
Lemma 4.4
Define $\hbar_{m}=(2m(\ln(m+1)-\ln\eta))^{1/2}$. In the event that $\theta_{k}>\theta^{*}$, $\mathbb{P}(Z_{k}(\theta_{k})=+1|\theta_{k})\leq\eta/2$. In the event that $\theta_{k}<\theta^{*}$, $\mathbb{P}(Z_{k}(\theta_{k})=-1|\theta_{k})\leq\eta/2$. That is, the detection accuracy $p(\theta)$ when the vertical move stops is bounded by $p(\theta)\geq 1-\eta/2$.
Lemma 4.4 states that the detection accuracy is lower bounded by $p_{c}:=1-\eta/2$, where the accuracy can be chosen through the selection of $\eta$. As $\theta$ approaches $\theta^{*}$, differentiating between choices becomes increasingly challenging because the drift tends to diminish, potentially requiring more steps in the testing process. However, our objective, you must recall, is to locate a $\theta$ such that $|\theta-\theta^{*}|\leq\varepsilon$, with a confidence level of $1-\delta$. In the vicinity of $\theta^{*}$, waiting until $|S_{m}(\theta)|\geq\hbar_{m}$ to terminate the test may not be necessary. A high volume of queries at a particular point $\theta_{k}$ suggests close proximity between $\theta_{k}$ and $\theta$, so that considering a earlier termination of the test may be advisable.
Based on this insight, we introduce a stopping time $\tau^{\uparrow}(\delta,\varepsilon)$, which is determined by the precision and confidence levels. This stopping time regulates the maximum number of queries carried out at a single point as $\theta$ approaches the local neighborhood of the true parameter (a concept we elaborate on in Section 4.4). Specifically, $\theta$ converges toward $\theta^{*}$ in a defined local neighborhood, denoted as $\{\theta^{\prime}:d(\theta,\theta^{\prime})\leq a\}$ for some positive constant $a$; here, surpassing the threshold $\tau^{\uparrow}(\delta,\varepsilon)$ signifies the difficulty in distinguishing between the parameters because of their proximity, and the algorithm terminates.
4.2.2 Horizontal moves
Algorithm 2 begins with a prior density $\rho_{0}$ on $[-\beta_{\Theta},\beta_{\Theta}]$ that is positive everywhere. For example, $\rho_{0}$ can be chosen as a uniform distribution. Choose some $p\in(1/2,p_{c})$, and let $q=1-p$. When the algorithm queries at the point $\theta_{k}$ to obtain $Z_{k}$, we update the density $\rho_{k}$:
  1. (I)
    If $Z_{k}(\theta_{k})=+1$, then
    $\rho_{k+1}(\theta^{\prime})=\begin{cases}2p\rho_{k}(\theta^{\prime}),&\mbox{if}\theta^{\prime}\geq\theta_{k};\\ 2q\rho_{k}(\theta^{\prime}),&\mbox{if }\theta^{\prime}<\theta_{k}.\end{cases}$
  2. (II)
    If $Z_{k}(\theta_{k})=-1$, then
    $\rho_{k+1}(\theta^{\prime})=\begin{cases}2q\rho_{k}(\theta^{\prime}),&\mbox{if}\theta^{\prime}\geq\theta_{k};\\ 2p\rho_{k}(\theta^{\prime}),&\mbox{if }\theta^{\prime}<\theta_{k}.\end{cases}$
The next query point $\theta_{k+1}$ is determined as the median of $\rho_{k+1}$. The probability density on the correct side would be increased while the probability density on the wrong side would be reduced. Define $\tau^{\rightarrow}(\delta,a)$ as the number of horizontal moves to reach $a$-neighborhood of $\theta^{*}$ with confidence level $1-\delta$ (which will be specifically defined in the paper later). That is, parameter $\theta$ moves to the local $a$-neighborhood of $\theta^{*}$ with probability of at least $21-\delta/$ 2 when the number of steps exceeds $\tau^{\rightarrow}(\delta/2,a)$. The horizontal move stops when the total number of moves reaches $\tau^{\rightarrow}(\delta/2,\varepsilon)$ because at this point the precision level $\varepsilon$ has been achieved.
In summary, there are two scenarios:
  1. (I)
    If the horizontal move is less than $\tau^{\rightarrow}(\delta/2,a)$ (indicating it may not yet reach the local neighborhood of $\theta^{*}$ within a distance of $a$), the vertical test stops when the movement reaches $\tau_{1}^{\uparrow}(\theta)$. Then it moves to the next query point $\theta$.
  2. (II)
    If the horizontal move is more than $\tau^{\rightarrow}(\delta/2,a)$, signifying entry into the local neighborhood of $\theta^{*}$ within a distance of $a$, there are two possible options:
    1. (a)
      When the vertical move first reaches $\tau^{\uparrow}(\delta/2,\varepsilon)$, the algorithm terminates, having identified a $\theta$ within the $\varepsilon$ range of $\theta^{*}$.
    2. (b)
      When the vertical move first reaches $\tau_{1}^{\uparrow}(\theta)$, the vertical test stops and then transitions to the next query point $\theta$.
The rationale behind distinguishing between local and non-local scenarios lies in the observation that two distinct situations can result in a significant number of vertical moves: (1) when both compared models are performing well (i.e., $\theta_{1}$ and $\theta_{2}$ are close to $\theta^{*}$), or (2) when both models are performing poorly (i.e., $\theta_{1}$ and $\theta_{2}$ are far from $\theta^{*}$). In the latter case, that the algorithm not terminate prematurely is crucial because $\theta$ has not yet gotten sufficiently close to $\theta^{*}$. Therefore, the stopping criterion $\tau^{\uparrow}(\delta,\varepsilon)$ is only applied when $\theta$ transitions into the local neighborhood of $\theta^{*}$.
Algorithm 2 is described as follows. We initialize the distribution as $\mu_{0}$. Without any historical data, $\mu_{0}$ can be initialized as the uniform distribution over the confidence region. Alternatively, $\mu_{0}$ can be approximated by the empirical distribution of the estimator using the bootstrap method and applying it to the first phase. At step $k$, the algorithm tests $\theta_{k}$ by collecting $Z_{k,s}(\theta_{k})$. If $\theta$ is located at the local neighborhood of $\theta^{*}$ ($k\geq\tau^{\rightarrow}(\delta/2,a)$) and the current step label $s$ is larger than $\tau^{\uparrow}(\delta/2,\varepsilon)$, then the algorithm stops and $\theta=\theta_{k}$. Otherwise, the vertical move stops when $s$ is larger than $\tau_{1}^{\uparrow}(\theta_{k})$. Once the test on $\theta_{k}$ stops, we update the prior distribution $\rho_{k}$ and select the next $\theta_{k+1}$ as the median of the updated distribution.
Algorithm 2 MAPB (Model Alignment through Probabilistic Bisection)
1:input: confidence level $\delta$, precision $\varepsilon$; initial distribution $\mu_{0}$; local parameter $a>0$; $k=0$; $\eta$;
2:repeat
3:     pick $\theta_{k}$ as the median of $\mu_{k}$
4:     $s=0$
5:     repeat
6:         test $\theta_{k}$ by collecting $Z_{k,s}(\theta_{k})$
7:         $s=s+1$
8:         if $k\geq\tau^{\rightarrow}(\delta/2,a)$ and step $s$ satisfies stopping criteria $\tau^{\uparrow}(\delta/2,\varepsilon)$ then
9:              Algorithm terminates and $\theta=\theta_{k}$
10:         end if
11:     until vertical step $s$ satisfies stopping criteria $\tau_{1}^{\uparrow}(\theta_{k})$
12:     $k=k+1$; update prior distribution $\rho_{k}$
13:until horizontal step $k$ satisfies stopping criteria $\tau^{\rightarrow}(\delta,\varepsilon)$
14:pick $\theta_{k}$ as the median of $\mu_{k}$
4.3 Complexity of Horizontal Moves
Let $T^{\uparrow}(\theta)$ denote the number of comparisons that are collected at $\theta$ and let $T^{\rightarrow}$ denote the number of horizontal moves; let $T_{k}^{+}$ denote the total number of comparisons that are collected after $k$ horizontal moves, where “$+$” denotes both horizontal and vertical moves. Note that $T^{\uparrow}(\theta)$, $T^{\rightarrow}$, and $T_{k}^{+}$ are all random variables. In the following paragraphs, we discuss the complexity of vertical and horizontal moves, and we characterize its sample complexity.
We first characterize the complexity of horizontal moves. Fix some $a\in(0,\theta^{*}+\beta_{\Theta})$. Define three regions: $A=[-\beta_{\Theta},\theta^{*}-a]$, $B=(\theta^{*}-a,\theta^{*}]$, and $C=(\theta^{*},\beta_{\Theta}]$. Define $T(a)=\inf\{k^{\prime}\geq 0:\theta_{k}\in B\cup C,\forall k\geq k^{\prime}\}$ to be the time required until the sequence of medians never reenters $A$. Let $\nu_{k}$ denote the (random) measure corresponding to the conditional posterior distribution, conditional on lying to the left of $\theta^{*}$, so that for any measurable $D$, $\nu_{k}(D)=\mu_{k}(D\cap[-\beta_{\Theta},\theta^{*}])/\mu_{k}([-\beta_{\Theta},\theta^{*}])$. Next, fix $\Delta\in(0,1/2)$ and define the stopping time $\tau(a)$ to be the first time that the conditional mass in $B$ lies above $1-\Delta$ and the median $\theta_{k}$ lies to the left of $\theta^{*}$; that is,
$\tau(a)=\inf\{k\geq 0:\nu_{k}(B)>1-\Delta,\theta_{k}<\theta^{*}\}.$
In line with the approach taken by Frazier et al. (2019) in establishing a connection between $T(a)$ and $\tau(a)$, we present a similar result in Lemma 4.5.
Lemma 4.5
It holds that
$T(a)\leq_{s.t.}\tau(a)+R,$ (4.1)
where $R$ is a random variable that is independent of $\tau(a)$ and $a$.
From (4.1), to bound $T(\cdot)$, we characterize the distribution of $\tau(\cdot)$. Algorithm 2 starts with an initial distribution $\mu_{0}$. Without any prior information, $\mu_{0}$ can be set as a uniform distribution; otherwise, we can use the prior information to update $\mu_{0}$. First, we make the following mild assumption regarding the initial prior distribution.
{assumption}
[Prior distribution] The initial prior distribution satisfies that $\mu_{0}(B)\geq a^{\varsigma}$ for some $\varsigma\leq 1$.
The uniform distribution satisfies Assumption 4.5 because $\mu_{0}(B)=\mu_{0}((\theta^{*}-a,\theta^{*}])=a$. With additional information, the prior distribution would be more centered around the true parameter $\theta^{*}$. Thus, $\varsigma$ is assumed to be less than or equal to 1. Under this assumption of prior distribution, we analyze the distribution of $\tau(\cdot)$ in Lemma 4.6.
Lemma 4.6
Under Assumption 4.5, by setting $r=r_{2}\alpha/(4\varsigma)$,
$\mathbb{P}(\tau(\omega e^{-rk})>k/2)\leq e^{-r_{2}\alpha k/4}\cdot\frac{\theta^{*}}{\omega^{\varsigma}}+\beta e^{-r_{1}k/2},$
for some $r_{1},r_{2},\alpha>0$ and any $\omega>0$.
Lemma 4.6 shows the light tail distribution of $\tau(\cdot)$. For small $\omega$, note that $\omega^{\varsigma}$ decreases in $\varsigma$. That is, when the initial prior distribution is centered more closely around the true parameter, $\varsigma$ is smaller; thus, the right-hand-side tail distribution is lighter.
We next characterize the complexity of horizontal moves.
Theorem 4.7 (Sample complexity of horizontal moves)
Under Assumption 4.5, by setting
$\tau^{\rightarrow}(\delta,\varepsilon)=\max\left\{\frac{4\varsigma(1+\varsigma)\log(\frac{8\theta^{*}}{\delta\varepsilon^{\varsigma}})}{r_{2}\alpha},\frac{2\log(8\beta_{2}/\delta)}{r_{1}},\frac{\log(8\beta_{2}/\delta)}{r_{R}}\right\}.$
we have that
$\mathbb{P}(|\theta_{k}-\theta^{*}|\leq\varepsilon)>1-\delta,\quad\forall k\geq\tau^{\rightarrow}(\delta,\varepsilon)\,.$
Theorem 4.7 states that after $\tau^{\rightarrow}(\delta,\varepsilon)$ horizontal moves, $\theta_{k}$ reaches the $\varepsilon$-neighborhood of the true parameter with probability at least $1-\delta$. The complexity of this process depends on $\log(1/\delta)$, $\log(1/\varepsilon)$, and $\varsigma(1+\varsigma)$. When the initial prior distribution is more closely centered around the true parameter $\theta^{*}$, fewer horizontal moves are needed.
Define the local region of $\theta^{*}$ as $[\theta^{*}-a,\theta^{*}+a]$. We divide the horizontal moves into two phases: Phase I includes all steps until the sequence of the medians never reenters $[-\beta_{\Theta},\theta^{*}-a]\cup[\theta^{*}+a,\beta_{\Theta}]$; and Phase II includes all steps after Phase I. Similar to the definition of $T(a)$, we define $T^{\prime}(a)$ as the stopping time, which is when the medians never reenter $A^{\prime}=[\theta^{*}+a,\beta_{\Theta}]$. Let $\psi(a)$ be the time required until the sequence of the medians never reenters $A\cup A^{\prime}$ – that is, $\psi(a)=\max\{T(a),T^{\prime}(a)\}$. Proposition 4.8 provides an upper bound for the expectation of moves outside of $[\theta^{*}-a,\theta^{*}+a]$.
Proposition 4.8
The expectation of $\psi(a)$ is bounded by
$\mathbb{E}[\psi(a)]\leq\frac{4\theta^{*}}{r_{2}\alpha a^{\varsigma}}+\frac{4\beta_{1}}{r_{1}}+\frac{2\beta_{2}}{r_{R}}.$
4.4 Complexity of Vertical Moves
The vertical move has two stopping criteria that depend on whether $\theta_{k}$ has moved to the local neighborhood of $\theta^{*}$. If $\theta_{k}$ is still outside of the local area, the algorithm would finish with the power-one test; otherwise, the algorithm may terminate earlier, before stopping the power-one test.
Lemma 4.9 characterizes the expected number of queries needed for testing $\theta$.
Lemma 4.9 (Power-one test)
For any $\theta\neq\theta^{*}$,
$\mathbb{E}[\tau_{1}^{\uparrow}(\theta)]\leq c_{1}|2\tilde{p}(\theta)-1|^{-2}\ln(|2\tilde{p}(\theta)-1|^{-1})+c_{2}\,,$
for some $c_{1},c_{2}>0$.
Lemma 4.9 implies that the expected number of vertical moves increases in the order of $O((2\tilde{p}(\theta)-1)^{-2})$ when $\tilde{p}(\theta)$ approaches 1/2. When $\tilde{p}(\theta)$ is bounded away from $1/2$, the expected vertical move is also bounded. As a result of Lemma 3.7, $\tilde{p}(\theta)$ is strictly larger than 1/2 for any $\theta\neq\theta^{*}$; thus, according to Heine–Borel theorem, there exists $\delta^{\prime}>0$ such that $\tilde{p}(\theta)>1/2+\delta^{\prime}$ for all $\theta\in[-\beta_{\Theta},\theta^{*}-a]\cup[\theta^{*}+a,-\beta_{\Theta}]$. (See the proof of Lemma 9.9 in Appendix 9.1.) For the ease of notation, we define the bound of vertical moves:
$\phi(\tilde{p})=c_{1}|2\tilde{p}-1|^{-2}\ln(|2\tilde{p}-1|^{-1})+c_{2}.$
The selection accuracy’s being bounded away from $1/2$ implies that $\max_{a\leq\|\theta-\theta^{*}\|\leq\beta_{\Theta}}\phi(\tilde{p}(\theta))$ also is bounded. Thus, we can infer the expected number of steps—comprising both horizontal and vertical movements—outside of $[\theta^{*}-a,\theta^{*}+a]$ as follows.
Proposition 4.10 (Sample complexity of Phase I)
The expected steps outside of $[\theta^{*}-a,\theta^{*}+a]$ is bounded by
$\mathbb{E}\left[\sum_{k=1}^{\psi(a)}T^{\uparrow}(\theta_{k})\boldsymbol{1}(\|\theta_{k}-\theta^{*}\|\geq a)\right]\leq\left(\frac{4\theta^{*}}{r_{2}\alpha a^{\varsigma}}+\frac{4\beta_{1}}{r_{1}}+\frac{2\beta_{2}}{r_{R}}\right)\max_{a\leq\|\theta-\theta^{*}\|\leq\beta_{\Theta}}\phi(\tilde{p}(\theta))\,.$
Proposition 4.10 highlights that the sample complexity of Phase I hinges on the choice of $a$ and the span of $\theta$ from the prior distribution. If $\beta_{\Theta}$ is overly expansive, there is a high probability that both selected parameters $\theta_{1}$ and $\theta_{2}$ are distant from the true parameter. Consequently, the accuracy of human selection may diminish considerably – and may even resemble random selection, as illustrated in Example 4.2. Thus, by narrowing the range of $\theta$ based on supplementary information, the sample complexity of Phase I can be notably diminished.
Once Phase I ends, $\theta_{k}$ moves to the $a$-neighborhood of $\theta^{*}$. In scenarios where both options are close to $\theta^{*}$, discerning the superior parameter becomes more challenging. Depending on whether the selection accuracy converges toward $1/2$ as the detected parameter approaches the true parameter, we classify the problem into two cases: 1) The selection accuracy is bounded away from 1/2 for all $\theta\neq\theta^{*}$; and 2) the section accuracy approaches $1/2$. The second case is harder and requires more delicate analysis. We first characterize the sample complexity for Case 1 and then analyze Case 2.
For the scenario where $\tilde{p}$ is bounded away from $\underline{p}>1/2$, combining the bound on the vertical moves (Proposition 4.10) with the complexity of the horizontal moves (Theorem 4.7), Theorem 4.11 characterizes the total sample complexity.
Theorem 4.11 (Sample complexity when $\underline{p}$ exists)
If $\tilde{p}(\theta)\geq\underline{p}>1/2$ for all $\theta\in\Theta$, then the expected number of samples collected before algorithm termination is
$\mathbb{E}[T^{+}_{T^{\rightarrow}}]\leq\tau^{\rightarrow}(\delta,\varepsilon)\cdot\phi(\underline{p})=O(\log(1/(\delta\varepsilon))).$
Next, we address the more challenging scenario, in which the selection accuracy tends toward $1/2$ as the tested parameter approaches the true parameter.
4.4.1 When the selection accuracy approaches $1/2$.
As the comparison accuracy approaches $1/2$, the vertical move (based on the power-one test) requires the collection of more samples. The rate at which the vertical sample complexity increases depends on the rate at which the utility changes as $\theta$ converges to $\theta^{*}$. To account for this effect, we introduce the following assumption.
{assumption}
For any $\theta$ such that $\|\theta-\theta^{*}\|\leq a$, there exists $\lambda_{\Delta}>0$ and $0<\kappa\leq 1$ such that
$|u(c_{\Delta}^{+}(\theta))-u(c_{\Delta}^{-}(\theta))|\geq\lambda_{\Delta}\|\theta-\theta^{*}\|^{\kappa}\,.$
The parameter $\kappa$ characterizes the speed of convergence, with a smaller value indicating a faster convergence rate. Assuming that $\kappa\leq 1$ is natural for several reasons. First, human decision making should not be inferior to learning from a random sample; otherwise, collecting random samples would suffice. We demonstrate this by examining the selection accuracy of learning from a single sample. Suppose the sample $(x,y)$ is generated from $y=\theta^{*}x+\epsilon$ where $\epsilon\sim\mathcal{N}(0,\sigma^{2})$. When comparing $\theta+\Delta$ and $\theta-\Delta$ (without loss of generality, assuming $\theta<\theta^{*}$), $\theta+\Delta$ is preferable to $\theta-\Delta$ if the residual is smaller. That is, the probability that $\theta+\Delta$ is better is
$\mathbb{P}(\theta+\Delta\succeq\theta-\Delta)$ $=\mathbb{P}\left((\theta^{*}x+\epsilon-(\theta+\Delta)x)^{2}\leq(\theta^{*}x+\epsilon-(\theta-\Delta)x)^{2}\right)$
$=\mathbb{P}(\epsilon\geq(\theta-\theta^{*})|x|)=\frac{1}{2}+\Omega((\theta^{*}-\theta)|x|).$
In instances where evaluation is conducted solely on a random sample, $\kappa$ is equal to 1. Considering that humans possess or have access to additional side information and thus can make better comparisons, we anticipate that $\kappa$ is, at most, 1.
Lemma 4.12 (How selection accuracy converges)
Under Assumption 4.4.1, it holds that
$\tilde{p}(\theta)-1/2\geq c\|\theta-\theta^{*}\|^{\kappa}\,,\forall\theta\in[\theta^{*}-a,\theta^{*}+a],$
where $c=\min\{\lambda_{\Delta}/(8\gamma\ln(2)),1/(6a^{\kappa})\}$.
Under Assumption 4.4.1, $\tilde{p}(\theta)$ converges to 1/2 at a speed no faster than the linear convergence. Therefore, when $\theta$ moves to the $a$-neighborhood of $\theta^{*}$, the distance between $\theta$ and $\theta^{*}$ can be bounded by $((\tilde{p}(\theta)-1/2)/c)^{1/\kappa}$. In other words, the closeness between $\tilde{p}(\theta)$ and 1/2 implies the closeness between $\theta$ and $\theta^{*}$. Thus, we can define $\tau^{\uparrow}(\delta,\varepsilon)$ as the threshold of the vertical test, such that if the vertical move $T^{\uparrow}(\theta)$ has reached $\tau^{\uparrow}(\delta,\varepsilon)$, then $\theta$ is close enough to $\theta^{*}$.
Theorem 4.13 (Vertical moves in Phase II)
Suppose that Assumption 4.4.1 holds. When $T^{\rightarrow}\geq\tau^{\rightarrow}(\delta/2,a)$ and $T^{\uparrow}(\theta_{T^{\rightarrow}})\geq\tau^{\uparrow}(\delta/2,\varepsilon)$, where
$\tau^{\uparrow}(\delta,\varepsilon)=\max\left\{\tau_{0},\frac{8\log(1/\delta)}{c^{2}\varepsilon^{2\kappa}}\right\}=O(\log(1/\delta)\varepsilon^{-2\kappa})\,,$
and $\tau_{0}=\max\{s:\hbar_{s}/s\geq c\varepsilon^{\kappa}/2\}$, where $c$ is defined in Lemma 4.12, then we have
$\mathbb{P}(\|\theta_{T^{\rightarrow}}-\theta^{*}\|\leq\varepsilon)\geq 1-\delta\,.$
When Assumption 4.4.1 holds and $\theta$ moves to the neighborhood of $\theta^{*}$, the sample complexity of the vertical move would not exceed $O(\varepsilon^{-2\kappa})$, where $\varepsilon$ is the precision. The vertical sample complexity increases as the value of $\kappa$ increases. The faster the selection accuracy converges to 1/2, the harder the selection problem, and thus the higher the vertical sample complexity. In the extreme case, where $\kappa=0$, $\tilde{p}(\theta)$ is bounded away from 1/2 so $\tau^{\uparrow}(\delta,\varepsilon)$ is bounded. This conclusion is consistent with what we have in Theorem 4.11.
4.5 Sample Complexity
Combining the sample complexity of horizontal moves and vertical moves, we conclude the total sample complexity of MAPB.
Theorem 4.14 (Sample complexity of MAPB)
Under Assumption 4.4.1, the expected number of moves is bounded by
$\mathbb{E}\left[\sum_{k=1}^{T^{\rightarrow}}T^{\uparrow}(\theta_{k})\right]\leq H(\delta,\varepsilon;a)=\tilde{O}(\varepsilon^{-2\kappa}),$
where
$H(\delta,\varepsilon;a)=\left(\frac{4\theta^{*}}{r_{2}\alpha a^{\varsigma}}+\frac{4\beta_{1}}{r_{1}}+\frac{2\beta_{2}}{r_{R}}\right)\max_{a\leq\|\theta-\theta^{*}\|\leq\beta_{\Theta}}\phi(\tilde{p}(\theta))+\tau^{\rightarrow}(\delta/2,\varepsilon)\tau^{\uparrow}(\delta/2,\varepsilon).$
Theorem 4.14 analyzes the (human) sample complexity in a one-dimensional space, which is in the order of $\tilde{O}(\varepsilon^{-2\kappa})$. The horizontal move is in the logarithm order, so the dominant part comes from the vertical moves in the neighborhood of $\theta^{*}$. If humans have the ability to differentiate between any two parameters $\theta_{1}$ and $\theta_{2}$, with the probability strictly higher than $1/2+\delta^{\prime}$ as long as $\theta_{1}\neq\theta_{2}$, then the total complexity is in the logarithm order of both confidence $1/\delta$ and precision $1/\varepsilon$.
In a $d$-dimensional space, if naively refining parameters along each dimension to make the parameter within an $\varepsilon$-ball (in $\ell_{\varrho}$-norm), then the total complexity is $dH(\delta/d,\varepsilon/d^{1/\varrho};a)$, which scales as $\tilde{O}(d^{1+\frac{2\kappa}{\varrho}}/\varepsilon^{2\kappa})$. However, in high-dimensional spaces, $d$ can be significant even though the true parameter may lie within a low-dimensional space. In the following section, we explore leveraging a supervised learning oracle to reduce sample complexity.
5 Model Alignment: Human-AI Interaction
In the previous section, we discussed the framework of using human feedback in a one-dimensional space. In practice, a noisy labeled dataset usually can be acquired easily. A noisy labeled dataset potentially can be used for learning the low-dimensional representation $\varphi(\cdot)$ and thus for reducing the complexity of the refining step. In this section, we propose a two-stage framework where the first stage learns the low-dimensional representation using noisy labels and the second stage refines the model using human feedback.
5.1 Two-stage Framework
Suppose the true parameter lies in an $s$-dimensional space where $s\ll d$. Given confidence $\delta$ and precision $\varepsilon$, we introduce a two-stage framework SL+LHF:
  1. (I)
    In Stage 1, the algorithm feeds $N_{1}^{HAI}(\delta,\varepsilon)$ noisy labeled data $(\boldsymbol{X},Y)$ to the supervised learning oracle (e.g., Lasso), where $\boldsymbol{X}\in\mathbb{R}^{N_{1}^{HAI}\times d}$ is the design matrix and $Y$ is the noisy response. At the end, we learn the $s$-dimensional embedding $\varphi(\cdot)$.
  2. (II)
    In Stage 2, the algorithm asks a human to refine the estimator along $s$ dimensions. This stage requires $N_{2}^{HAI}(\delta,\varepsilon)$ data points.
Algorithm 3 SL+LHF (Supervised Fine-Tuning+Human Comparison)
1:Input: Confidence level $\delta$, precision $\varepsilon$; initial distribution $\mu_{0}$; local parameter $a>0$; $k=0$; $\eta$;
2:Stage 1: Learn the embedding $\varphi(\cdot)$ through SL oracle trained on data $(\boldsymbol{X},Y)$ with size $N_{1}^{HAI}(\delta,\varepsilon)$;
3:Stage 2: Execute Algorithm 2 (MAPB) by querying $N_{2}^{HAI}(\delta,\varepsilon)$ samples.
The framework design is grounded in the insight that certain representation learning techniques, like autoencoders or deep learning architectures, are inherently robust to noisy labels (Li et al. 2021, Taghanaki et al. 2021). Stage 1 is dedicated to uncovering underlying patterns and structures within the data. Given the effectiveness of human expertise in model refinement, Stage 2 serves to enhance model alignment through targeted sample querying. In what follows, we specifically discuss sparse linear models (Example 3.1), where Stage 1 applies Lasso [^2] OLS and thresholding may serve the same purpose as Lasso. to select important features.
5.2 Illustration on Linear Models
To demonstrate the sample complexity reduction for the two-stage framework, compared to a framework trained purely by supervised learning, and to provide theoretical justifications, we focus on sparse linear models:
$Y_{i}=\left\langle\boldsymbol{\vartheta^{*}},X_{i}\right\rangle+\epsilon_{i}=\left\langle\boldsymbol{\theta}^{*},\varphi(X_{i})\right\rangle+\epsilon_{i},$
where $\varphi(\cdot)$ projects $X_{i}$ to all important features; $S$ is the index set of all important features; and the parameter $\boldsymbol{\theta}^{*}$ satisfies that $|\theta^{*}_{i}|\geq\underline{\beta}_{\Theta}>0$ for all $i\in S$. In the initial stage, we use Lasso for the selection of significant features, followed by human comparisons for model alignment. To assess the sample complexity of our two-stage framework, we sequentially quantify the size of samples needed for the two stages. First, we establish the following standard assumptions for the analysis of Lasso.
{assumption}
[sub-Gaussian noise] The observational noise is zero-mean i.i.d. sub-Gaussian with parameter $\sigma$.
{assumption}
For the random design matrix $\boldsymbol{X}\in\mathbb{R}^{n\times d}$ input in the first stage, we assume that the following three conditions hold:
  1. (I)
    Mutual incoherence: $\max_{j\in S^{c}}\ \|(\boldsymbol{X}_{S}^{\top}\boldsymbol{X}_{S})^{-1}\boldsymbol{X}_{S}^{\top}x_{j}\|_{1}\leq\alpha_{1};$
  2. (II)
    Lower eigenvalue: $\lambda_{\min}\left(V_{S}\right)\geq c_{\min}>0,$ where $V_{S}=\frac{1}{n}\boldsymbol{X}_{S}^{\top}\boldsymbol{X}_{S}$; and
  3. (III)
    $\ell_{\infty}$-curvature condition: $\|Vz\|_{\infty}\geq\alpha_{2}\|z\|_{\infty}\,,\text{ for all }z\in C_{\alpha^{\prime}}(S),$ where $C_{\alpha^{\prime}}(S)=\{\Delta\in\mathbb{R}^{d}|\|\Delta_{S^{c}}\|_{1}\leq\alpha^{\prime}\|\Delta_{S}\|_{1}\}$.
Assumption 5.2 is standard for the analysis of Lasso. For example, consider a random design matrix $\boldsymbol{X}\in\mathbb{R}^{n\times d}$ with i.i.d. $\mathcal{N}(0,1)$ entries; we can easily verify that it satisfies mutual incoherence, lower eigenvalue, and $\ell_{\infty}$-curvature condition. Under Assumptions 5.2 and 5.2 and based on Corollary 7.22 in Wainwright (2019), we derive the following theorem that bounds the infinite distance between vector $\hat{\theta}_{S}$ and $\theta^{*}_{S}$.
Theorem 5.1
Under Assumptions 5.2 and 5.2, suppose that we solve the Lagrangian Lasso with regularization parameter
$\lambda_{n}=\frac{2C\sigma}{1-\alpha_{1}}\left\{\sqrt{\frac{2\log(d-s)}{n}}+\zeta\right\}$
for $\zeta=\frac{1}{4}\min\left\{\frac{\sqrt{c_{\min}}}{\sigma\underline{\beta}_{\Theta}},\frac{\alpha_{2}(1-\alpha_{1})}{2C\sigma\underline{\beta}_{\Theta}}\right\}$. Then, the optimal solution $\hat{\theta}$ is unique, with its support contained within $S$, and it satisfies the $\ell_{\infty}$-error bound
$\|\hat{\theta}_{S}-\theta^{*}_{S}\|_{\infty}\leq\underbrace{\frac{\sigma}{\sqrt{c_{\min}}}\left\{\sqrt{\frac{2\log s}{n}}+\zeta\right\}+\frac{1}{\alpha_{2}}\lambda_{n}}_{B_{n}}\,,$
all with probability at least $1-4e^{-\frac{n\zeta^{2}}{2}}$. Moreover, there is no false inclusion: The solution has its support set $\hat{S}$ contained within the true support set $S$.
Remark 5.2
Lasso is not limited to the linear function but can also be applied to the nonlinear function class (Plan and Vershynin 2016) or to a nonparametric variable selection (Li et al. 2023). Moreover, there is a rich stream of literature on representation learning that can be applied in the first stage (Bengio et al. 2013).
Theorem 5.1 implies that as long as $B_{n}\leq\underline{\beta}_{\Theta}$, the variable selection is consistent. Because $\zeta=\frac{1}{4}\min\left\{\frac{\sqrt{c_{\min}}}{\sigma\underline{\beta}_{\Theta}},\frac{\alpha_{2}(1-\alpha_{1})}{2C\sigma\underline{\beta}_{\Theta}}\right\}$, we only need
$\frac{\sigma}{\sqrt{c_{\min}}}\sqrt{\frac{2\log s}{n}}+\frac{1}{\alpha_{2}}\cdot\frac{2C\sigma}{1-\alpha_{1}}\sqrt{\frac{2\log(d-s)}{n}}\leq\frac{1}{2}\underline{\beta}_{\Theta}.$
Thus, if there are $n$ noisy samples in which
$n\geq\max\left\{\frac{32\sigma^{2}\log s}{\underline{\beta}_{\Theta}^{2}c_{\min}},\frac{128C^{2}\sigma^{2}\log(d-s)}{(\underline{\beta}_{\Theta}\alpha_{2}(1-\alpha_{1}))^{2}}\right\},$
we are able to select $s$ variables with probability at least $1-4e^{-\frac{n\zeta^{2}}{2}}$. Define
$H_{0}(\delta,\varepsilon;\underline{\beta}_{\Theta})=\max\left\{\frac{32\sigma^{2}\log s}{\underline{\beta}_{\Theta}^{2}c_{\min}},\frac{128C^{2}\sigma^{2}\log(d-s)}{(\underline{\beta}_{\Theta}\alpha_{2}(1-\alpha_{1}))^{2}},\frac{2\log(4/\delta)}{\zeta^{2}}\right\}=O(\sigma^{2}\log d/\underline{\beta}_{\Theta}^{2}).$
Then, when we have more than $H_{0}(\delta,\varepsilon;\underline{\beta}_{\Theta})$ noisy labeled data at hand, we can have the correct feature selection with probability at least $1-\delta$. Moreover, if we want to undertake one more step of refinement to restrict the uncertainty set to be within $\mathfrak{r}$-distance, where $\mathfrak{r}<\underline{\beta}_{\Theta}$, then we need $H_{0}(\delta,\varepsilon;\mathfrak{r})$ noisy labeled data. Note that $\mathfrak{r}$, which represents the feasible range for refinement, is the parameter that we can choose. The trade-off exists: A smaller value of $\mathfrak{r}$ demands more noisy labels in the initial stage. However, limiting the feasible region to a smaller area would conserve the samples that are required for human comparisons. As previously noted, when both options significantly deviate from the true model, distinguishing the superior one becomes challenging. This challenge implies a need for more samples to ascertain the correct answer. Consequently, in certain scenarios, opting for $\mathfrak{r}\leq\underline{\beta}_{\Theta}$ to confine the refinement region may be optimal. Specifically, we select $\mathfrak{r}$ to minimize the total sample complexity:
$\mathfrak{r}^{*}=\min_{0<\mathfrak{r}\leq\underline{\beta}_{\Theta}}H_{0}(\delta,\varepsilon;\mathfrak{r})+sH(\delta/s,\varepsilon/s^{1/\varrho};\mathfrak{r}),$ (5.1)
where $H(\delta,\varepsilon;\mathfrak{r})$ is defined in Theorem 4.14 and we measure the distance using $l_{\varrho}$-norm. Note that although the first part depends on the noise $\sigma$, the second part relies not on $\sigma$ but on the detection accuracy $\phi(\tilde{p}(\theta))$. This point is important to emphasize. Hence, the optimal $\mathfrak{r}^{*}$ is contingent on the function of detection accuracy. For $\theta$ considerably distant from the true parameter $\theta^{*}$, $\tilde{p}(\theta)$ could potentially approach $1/2$, leading to a substantial value for $\phi(\tilde{p}(\theta))$.
5.3 Complexity of SL+LHF
The optimal value of $\mathfrak{r}^{*}$ can be obtained numerically by enumerating and then comparing the objectives in Equation (5.1). To establish an upper bound on the sample complexity for the two-stage framework so that we analytically show its benefit, we set $\mathfrak{r}=\underline{\beta}_{\Theta}$. In this case, we have
$H_{0}(\delta,\varepsilon;\underline{\beta}_{\Theta})=\tilde{O}(\sigma^{2}/\underline{\beta}_{\Theta}^{2})\quad\text{and}\quad sH(\delta/s,\varepsilon/s^{1/\varrho};\underline{\beta}_{\Theta})=\tilde{O}\left(s\bar{\phi}(a,\underline{\beta}_{\Theta})+s^{1+\frac{2\kappa}{\varrho}}/\varepsilon^{2\kappa}\right),$
where $\bar{\phi}(a,D)=\max_{a\leq\|\theta-\theta^{*}\|\leq\underline{\beta}_{\Theta}}\phi(\tilde{p}(\theta))$. Theorem 5.3 is a direct conclusion from Theorems 4.14 and 5.1.
Theorem 5.3 (Computational complexity of SL+LHF)
Under Assumptions 5.2 and 5.2, set $N_{1}^{HAI}(\delta)=H_{0}(\delta,\varepsilon;\underline{\beta}_{\Theta})$ and then run Algorithm 2 with parameter range $[-\underline{\beta}_{\Theta},\underline{\beta}_{\Theta}]$; then, it holds that $\|\hat{\theta}-\theta^{*}\|_{\varrho}\leq\varepsilon$ with probability at least $1-2\delta$. Moreover, under Assumption 4.4.1, the expected sample complexity is
$N_{1}^{HAI}(\delta,\varepsilon)+N_{2}^{HAI}(\delta,\varepsilon)=\tilde{O}(\sigma^{2}/\underline{\beta}_{\Theta}^{2}+s\bar{\phi}(a,\underline{\beta}_{\Theta})+s^{1+\frac{2\kappa}{\varrho}}/\varepsilon^{2\kappa}).$
Theorem 5.3 demonstrates the sample complexity of the two-stage framework SL+LHF. When the distance is measured in two-norm (i.e., $\varrho=2$), the complexity is in the order of $\tilde{O}(\sigma^{2}/\underline{\beta}_{\Theta}^{2}+s\bar{\phi}(a,\underline{\beta}_{\Theta})+s^{1+\kappa}/\varepsilon^{2\kappa})$. Instead, if Lasso is used exclusively for the supervised learning oracle to refine the estimator based on $n$ noisy labeled data (subject to certain regularity conditions), the estimation error can be bounded as $\|\hat{\theta}_{n}-\theta^{*}\|_{2}\leq c^{\prime}\sigma\sqrt{s\log d/n}$ with high probability for some constant $c^{\prime}>0$. Equivalently, we need to acquire $\tilde{O}(\sigma^{2}s\log d/\varepsilon^{2})$ data points to ensure that the two-norm distance of the estimation error is within $\varepsilon$.
To compare the sample complexity of SL+LHF as indicated by Theorem 5.3 for $\varrho=2$, we characterize the condition of SL+LHF that requires less sample complexity than SL in Proposition 5.4.
Proposition 5.4
For sufficiently small $\varepsilon$, SL+LHF requires less sample complexity than pure SL for solving the $(\varepsilon,\delta)$-alignment problem (ignoring the logarithm term) when:
$\frac{\sigma^{2}}{s^{\kappa}}\gtrapprox\varepsilon^{2-2\kappa}.$ (5.2)
We call the ratio $\sigma^{2}/s^{\kappa}$ the label noise-to-comparison-accuracy ratio (LNCA ratio), and we call Condition (5.2) the LNCA condition. This condition involves two important parameters: the observational noise $\sigma$ and the convergence rate of the accuracy $\kappa$. When the comparison accuracy is bounded away from 1/2, we have $\kappa=0$, and the LNCA condition reduces to $\sigma\gtrapprox\varepsilon$, which naturally holds true. In the worst case scenario, where $\kappa=1$, the LNCA condition becomes $\sigma^{2}\gtrapprox s$. When the observational noise of the sample is higher, the benefit of the human comparison step becomes more apparent.
As previously mentioned, our framework can be extended to scenarios where the tasks differ between the initial feature-learning stage and the final implementation stage, provided the low-dimensional representation remains consistent. This is because the sample complexity of the second refinement stage in our algorithm remains unaffected as long as the representation stays unchanged. Meanwhile, Lasso must relearn the varying parameters. Therefore, the sufficient condition outlined in Proposition 5.4 continues to hold.
6 Practical Implementations
In the preceding sections, we established a framework that assumes the opportunity for humans to compare any two parametric models. However, such a comparison could be challenging when humans have access only to the models themselves. In practice, comparing models based on their performance with sampled data is often more useful because this engagement allows humans to discern which model aligns better with the true underlying model. To address this specific challenge, in this section, we present the practical aspects of comparing two models using selected samples.
6.1 Active Learning: How to Select Samples for Comparison
Active learning focuses on the strategic selection of samples to be labeled or compared in our context, with the aim of constructing prediction models in a resource-efficient manner. The key idea is to assess the significance of a sample to determine the value of acquiring its labels. Algorithm 2 actively selects a pair of models for comparison. In practical scenarios, this choice often involves soliciting human judgment to compare the responses predicted by two different models for the given sample data, rather than directly comparing their model parameters.
We start with an example. The clinician needs to evaluate the atherosclerotic cardiovascular disease (ASCVD) risk of patients $f_{\boldsymbol{\theta}}$, which is a linear function of contextual information, denoted as $x$. Such information comprises demographics (age, gender, racial/ethnic group, and education level); baseline conditions (body mass index, history of ASCVD events, and smoking status); clinical variables (glycosylated hemoglobin (A1c) and systolic blood pressure (sbp)); anti-hypertensive agents; and anti-hyperglycemic agents.
Given a sample of patients’ contexts $x\in\mathcal{X}$, clinicians evaluate the ASCVD risk according to their expertise. The distance between models $f_{\boldsymbol{\theta}}$ and $f_{\boldsymbol{\theta}^{*}}$, evaluated on sample $x$, can be defined as the absolute value of the difference:
$d_{x}(\boldsymbol{\theta},\boldsymbol{\theta}^{*})=|\varphi(x)^{\top}\boldsymbol{\theta}-\varphi(x)^{\top}\boldsymbol{\theta}^{*}|.$
Then, the utility of $f_{\boldsymbol{\theta}}$ on sample $x$ is $u_{x}(\boldsymbol{\theta})=-d(\boldsymbol{\theta},\boldsymbol{\theta}^{*}|x)=-|\varphi(x)^{\top}\boldsymbol{\theta}-\varphi(x)^{\top}\boldsymbol{\theta}^{*}|$. When presented with risk levels predicted by $f_{\boldsymbol{\theta}_{1}}$ and $f_{\boldsymbol{\theta}_{2}}$, clinicians would typically opt for the model that produces results closest to their assessment.
Let us first consider a simple scenario. When $\mathcal{X}$ contains an adequate variety of samples, such that for any base vector $e_{i}$, there exists sample $x$ such that $\varphi(x)=e_{i}$, then by comparing $f_{\boldsymbol{\theta}_{1}}$ with $f_{\boldsymbol{\theta}_{2}}$ on this sample $x$, we have
$d_{x}(\boldsymbol{\theta}_{1},\boldsymbol{\theta}^{*})=|\theta_{1i}-\theta_{i}^{*}|\quad\text{and}\quad d_{x}(\boldsymbol{\theta}_{2},\boldsymbol{\theta}^{*})=|\theta_{2i}-\theta_{i}^{*}|.$
It implies that the comparison of $\boldsymbol{\theta}_{1}$ and $\boldsymbol{\theta}_{2}$, when testing on $x$, would focus only on the $i^{th}$ dimension. Thus, the alignment process for the parameter in different dimensions can be decoupled, and SL+LHF can be directly applied, where the Stage 2 alignment proceeds along each dimension.
However, in practice, the dataset $\mathcal{X}$ may be limited, potentially leading to situations where there are no samples $x\in\mathcal{X}$ that satisfy the condition $\varphi(x)=e_{i}$. In such cases, refining the model parameter along each dimension separatelybecomes impractical. The question then arises: How can we efficiently select samples to refine the model parameter? To address this challenge, we propose a procedure for constructing a new basis of comparison using the available samples in $\mathcal{X}$. This procedure aims to optimize the alignment process by leveraging the existing data to guide the selection of samples for parameter enhancement.
6.2 New Basis Construction through the Gram Schmidt Process
To facilitate the independent alignment of each coordinate, we aim to identify a new set of orthogonal basis, which can be constructed using $\Psi:=\{\varphi(x):x\in\mathcal{X}\}$. Let $\boldsymbol{\alpha}_{1},\cdots,\boldsymbol{\alpha}_{s}$ represent these new bases, and our objective is to learn $\boldsymbol{\theta}^{*}=\sum_{i=1}^{s}\omega_{i}^{*}\boldsymbol{\alpha}_{i}$, where $\omega_{i}^{*}$ are the corresponding coefficients. We describe the process of constructing the orthogonal basis $\{\boldsymbol{\alpha}_{1},\cdots,\boldsymbol{\alpha}_{s}\}$ from the sample space.
First, we pick a set of samples that spans the space of $\Psi$, denoted as $\varphi(x_{1}),\cdots,\varphi(x_{s})$. For simplicity, we define $\boldsymbol{z}_{i}=\varphi(x_{i})$. Then, we use the Gram Schmidt process to construct a set of orthogonal bases, based on $\boldsymbol{z}_{1},\cdots,\boldsymbol{z}_{s}$:
$\boldsymbol{\alpha}_{k}=\boldsymbol{z}_{k}-\sum_{j=1}^{k-1}\frac{\boldsymbol{z}_{k}^{\top}\boldsymbol{\alpha}_{j}}{\boldsymbol{\alpha}_{j}^{\top}\boldsymbol{\alpha}_{j}}\boldsymbol{\alpha}_{j},\quad\forall 1\leq k\leq s.$
We sequentially align the coordinates as $\omega_{1},\cdots,\omega_{s}$. When evaluating a model on sample $x_{1}$, the true value is represented by $\varphi(x_{1})^{\top}\boldsymbol{\theta}^{*}=\boldsymbol{z}_{1}^{\top}\boldsymbol{\theta}^{*}=\omega_{1}^{*}\boldsymbol{z}_{1}^{\top}\boldsymbol{z}_{1}$. Using a probabilistic bisection process in Algorithm 2, we actively select two answers for each query. Through this iterative refinement, the final value terminates at $y(x_{1})$, which is guaranteed to be within $\varepsilon$-accuracy of the true parameter $\omega_{1}^{*}$ (with high probability). In the end, the first coordinate is refined to $\hat{\omega}_{1}=y(x_{1})/(\boldsymbol{z}_{1}^{\top}\boldsymbol{z}_{1})$.
Next, the algorithm applies testing to the sample $\boldsymbol{z}_{2}=\varphi(x_{2})=\frac{\boldsymbol{z}_{2}^{\top}\boldsymbol{\alpha}_{1}}{\boldsymbol{\alpha}_{1}^{\top}\boldsymbol{\alpha}_{1}}\boldsymbol{\alpha}_{1}+\boldsymbol{\alpha}_{2}$. Note that the true value is given by
$\varphi(x_{2})^{\top}\boldsymbol{\theta}^{*}=\boldsymbol{z}_{2}^{\top}\boldsymbol{\theta}^{*}=\boldsymbol{z}_{2}^{\top}\boldsymbol{\alpha}_{1}\omega_{1}^{*}+\boldsymbol{\alpha}_{2}^{\top}\boldsymbol{\alpha}_{2}\omega_{2}^{*},$
where $w_{1}$ is already refined. Once more, using the probabilistic bisection process, the value of $z_{2}^{\top}\boldsymbol{\theta}^{*}$ is iteratively refined to $y(x_{2})$. Upon termination of the probabilistic bisection process, we obtain the aligned value of the second coordinate as
$\hat{\omega}_{2}=\frac{y(x_{2})-\boldsymbol{z}_{2}^{\top}\boldsymbol{\alpha}_{1}\hat{\omega}_{1}}{\boldsymbol{\alpha}_{2}^{\top}\boldsymbol{\alpha}_{2}}.$
Subsequently, at the $k^{th}$ iteration, we test on the sample $\boldsymbol{z}_{k}=\boldsymbol{\alpha}_{k}+\sum_{j=1}^{k-1}\frac{\boldsymbol{z}_{k}^{\top}\boldsymbol{\alpha}_{j}}{\boldsymbol{\alpha}_{j}^{\top}\boldsymbol{\alpha}_{j}}\boldsymbol{\alpha}_{j}$, where the true value equals
$\boldsymbol{z}_{k}^{\top}\boldsymbol{\theta}^{*}=\sum_{j=1}^{k-1}\omega_{j}^{*}\boldsymbol{z}_{k}^{\top}\boldsymbol{\alpha}_{j}+\omega_{k}^{*}\boldsymbol{\alpha}_{k}^{\top}\boldsymbol{\alpha}_{k}.$
After the value of $\boldsymbol{z}_{k}^{\top}\boldsymbol{\theta}^{*}$ is refined to $y(x_{k})$ through Algorithm 2, we can obtain the aligned value of the $k^{th}$ coordinate as
$\hat{\omega}_{k}=\frac{y(x_{k})-\sum_{j=1}^{k-1}\hat{\omega}_{j}\boldsymbol{z}_{k}^{\top}\boldsymbol{\alpha}_{j}}{\boldsymbol{\alpha}_{k}^{\top}\boldsymbol{\alpha}_{k}}.$
By asking a human to evaluate $x_{1},\cdots,x_{s}$ and by refining $\omega_{1},\cdots,\omega_{s}$ in sequence, we can refine the parameter $\boldsymbol{\theta}:=\sum_{i=1}^{s}\hat{\omega}_{i}\boldsymbol{\alpha}_{i}$ within a certain accuracy level. We include the pseudo code in Algorithm 4.
Note that the set of orthogonal bases is not unique; multiple sets may exist, and recognizing this possibility is essential. Consequently, we can conduct multiple rounds of alignment, with the option to switch the basis from one round to another.
Algorithm 4 ASS (Active Sample Selection)
1:input: available set of covariate $\mathcal{X}$
2:Pick a set of samples that spans the space of $\Psi$: $\varphi(x_{1}),\cdots,\varphi(x_{s})$
3:for $k=1,\cdots,s$ do
$\boldsymbol{\alpha}_{k}=\varphi(x_{k})-\sum_{j=1}^{k-1}\frac{\varphi(x_{k})^{\top}\boldsymbol{\alpha}_{j}}{\boldsymbol{\alpha}_{j}^{\top}\boldsymbol{\alpha}_{j}}\boldsymbol{\alpha}_{j}$
4:end for
5:for $k=1,\cdots,s$ do
6:     Query sample $x_{k}$ through Algorithm 2 and get output $y(x_{k})$
7:     Set $\hat{\omega}_{k}=\frac{y(x_{k})-\sum_{j=1}^{k-1}\hat{\omega}_{j}\varphi(x_{k})^{\top}\boldsymbol{\alpha}_{j}}{\boldsymbol{\alpha}_{k}^{\top}\boldsymbol{\alpha}_{k}}$
8:end for
7 Numerical Experiment
7.1 Synthetic Data Experiments
In the first experiment, we compare the performance of SL+LHF and SL using the same sample size to assess how parameter variations affect the performance of the two algorithms.
Experiment setup. Consider a high-dimensional linear regression problem where $d=100$ and $s=10$. The first ten coordinates of the true parameter are independently drawn from the uniform distribution, and the rest of the coordinates are zero. The observational noise is drawn from Gaussian distribution with a mean of zero and variance $\sigma\in\{1,2,5\}$; we choose $\kappa\in[0,0.95]$, where $\kappa$ comes from Assumption 4.4.1. Specifically, the difference of the utility in the two answers $c_{\Delta}^{+}(\theta)$ and $c_{\Delta}^{-}(\theta)$ is assumed to be $\|\theta-\theta^{*}\|^{\kappa}$. The selection precision (along each dimension) is set at $\varepsilon=0.1$, and the human expertise level in the choice model is $\gamma=1$. We conducted 30 repetitions for each experiment.
In the first stage of the two-stage framework, we selected a dataset with a sample size proportional to $\sigma^{2}\log d$ to input into Lasso. According to Theorem 5.1, this selection ensures that important features are identified with high probability. We chose the regularization parameter using cross-validation. The second stage is then executed, requiring human comparison samples, to reach the desired precision threshold $\varepsilon$. Once the SL+LHF algorithm completes, the combined sample size for both stages is recorded as the sample size for SL+LHF. Then, for a fair comparison, we provided the same sample size to SL (to train a lasso with the same number of observations) and computed the estimation error. We then looked at the ratio of the estimation errors of the two approaches, denoted by Err(SL+LHF)/Err(SL).
7.1.1 Experiment results.
Figure 3 shows the median of the ratio Err(SL+LHF)/Err(SL) using different values of $\kappa$ when $\sigma$ takes values $1$, $2$, or $5$. For all values of $\sigma$, when $\kappa$ is smaller, the human’s selection accuracy is higher as $\theta$ approaches $\theta^{*}$. In particular, when $\kappa=0$, the selection accuracy is bounded away from 50%. Therefore, the task is harder as $\kappa$ increases. It is consistent with the increasing trend in the estimation error ratio as $\kappa$ grows from 0 to 1. In addition, a larger value of $\kappa$ leads to a larger sample size, which leads to a decrease in the estimation error of lasso. When $\sigma=1$, the intersection point of the curve, with the horizontal line of the ratio equal to $1$ (red dashed line), is around $\kappa=0.2$. That is, when $\kappa\leq 0.2$, SL+LHF outperforms pure SL in more than half of the cases. In particular, when $\kappa=0$, the median of the error ratio is close to 0.
Figure 3: Error ratio of the two-stage framework to the pure SL with different values of $\sigma$.
The results for noise levels with $\sigma=2$ and $\sigma=5$ are also illustrated in Figure 3. Notably, the trends in the error curves are similar. The threshold, defined as the intersection point with the red dashed line, is approximately 0.4 when $\sigma=2$ and approximately $0.65$ when $\sigma=5$.
Overall, the three curves show that the superiority of SL+LHF becomes more pronounced in the presence of higher observational noise. In conclusion, SL+LHF is more effective under increased noise or reduced $\kappa$.
7.1.2 The variation of expertise level $\gamma$.
We vary the parameter $\gamma$ to model different levels of human expertise and to analyze how the accuracy of human comparisons affects the estimation error. For the experiment setup, we set $\sigma=2$ and $s=10$. The smaller value of $\gamma$ indicates higher accuracy in the comparison task. In particular, as $\gamma$ approaches 0, the human can always select the accurate answer. Figure 4 shows the performance of the two algorithms with $\gamma$ equal to $0.8$, $1$, and $1.1$. Comparing the three curves, when $\gamma=1.1$, SL+LHF performs better than pure SL if $\kappa$ is lower than approximately 0.3; when $\gamma=0.8$, SL+LHF performs better than pure SL if $\kappa$ is lower than approximately 0.5. Therefore, we find support for our theory that the performance of SL+LHF improves when the human has a higher expertise level in the selection (i.e., the selection accuracy is higher).
Figure 4: Error ratio of the two-stage framework to the pure SL with different values of $\gamma$.
7.1.3 The variation of important feature dimensions $s$.
The relative performance of two algorithms also depends on the dimension of the problem. For the experiment setup, we set $\gamma=1$ and $\sigma=2$. Figure 5 illustrates the algorithmic performance when important feature dimension $s$ takes values $10$, $20$, and $50$. The figure reveals that the intersection point shifts to a smaller value as the important feature dimension value increases. This observation aligns with our theoretical condition (5.2), indicating that as the value of the important feature dimension grows, the advantage of SL+LHF diminishes. This result highlights the importance of low-dimensional representations in the success of this method.
Figure 5: Error ratio of the two-stage framework to the pure SL with different values of $s$.
7.2 MTurk Experiment on the LNCA Ratio
In this section, we first outline the setup of our MTurk experiment [^3] https://www.mturk.com/. We then describe the estimation process for key parameters, including the accuracy convergence rate $\kappa$, expertise level $\gamma$, and noise level $\sigma$. After these parameters were estimated, we conducted simulations to compare the performance of the two-stage framework with that of the pure SL approach.
We designed a task in which participants are asked to estimate the number of points within a square. We divided participants into two distinct groups. The first group was presented with one of the squares and two options for the estimated number of points. The second group was shown the same square and asked to provide their estimated number of points. (See Figures 6 and 7.) The task for the second group was to select the option they believed was closer to the actual number. For instance, in Figure 6(b), the true number is 32. Human participants were provided with two options: 55 and 35; the correct choice is 35.
With this dual-group approach, we aimed to explore the efficacy of human estimation when participants were asked to estimate the number of points versus to choose from comparative options. In the latter case, we varied the discrepancy between the two choices to assess how the ease or difficulty of comparison affected the accuracy of participants’ selections. A larger disparity between the choices facilitated easier decision-making, while closer options posed a more challenging comparative task. Figure 6 shows three testing samples. Notably, sample (c) presents the most challenging comparison because of the minimal disparity ($|u(c^{+})-u(c^{-})|)$ observed between them. [^4] The gap between the first choice (45) and the true value (36) is 9 while the gap between the second choice (25) and the true value is 11. Thus, the difference between 9 and 11 is only 2.
To compare the performance of the two-stage framework and the pure supervised learning, we first estimated $\kappa$ and $\sigma$ from the online experiment in Amazon MTurk. Then we ran the simulation to compare two methods under different precision levels. We launched 50 different tasks, all structured in the same way as the task shown in Figure 7, and collected 500 data points in total. We estimated $\kappa$ and $\sigma$ as described in the following.
(a) 96 vs. 76. True: 64.
(b) 55 vs. 35. True: 32.
(c) 45 vs. 25. True: 36.
Figure 6: Experiment design for Group 1.
(a) Estimate the number of points.
Figure 7: Experiment design for Group 2.
Estimation process for $\kappa$. Given figures with true parameter value $\theta^{*}$ (i.e., the number of points in the figure), we estimate the accuracy convergence rate $\kappa$ between two options using a human choice model. We collect multiple samples for each figure and use these estimates to derive the relationship between the difference in utility and the accuracy of human choices.
Specifically, among two choices, assume the smaller one is $c^{-}$, the larger one is $c^{+}$, and the middle point is $\theta=(c^{-}+c^{+})/2$. According to the choice model, the human makes the correct choice with accuracy $p(\theta)=\frac{1}{1+\exp(-|u(c^{-})-u(c^{+})|/\gamma)}$. Note that we try to estimate $\kappa$ from Assumption 4.4.1 that
$|u(c^{+})-u(c^{-})|/\gamma\approx\lambda_{\Delta}/\gamma\cdot\|\theta-\theta^{*}\|^{\kappa}\,.$
That is, the accuracy is $p(\theta)=\frac{1}{1+\exp(-\lambda_{\Delta}/\gamma\cdot|\theta-\theta^{*}|^{\kappa})}$. Let $\tilde{\lambda}=\lambda_{\Delta}/\gamma$. For each data point with tested point $\theta_{i}$ and true parameter $\theta_{i}^{*}$, we define $y_{i}=1$ (the correct choice) with probability $p_{i}(\theta)$, and we define $y_{i}=0$ (the wrong choice) with probability $1-p_{i}(\theta)$. By maximizing the log-likelihood, we solve the optimization problem
$\max_{\tilde{\lambda},\kappa}\quad\sum_{i=1}^{K}y_{i}\log p_{i}+(1-y_{i})\log(1-p_{i}),$
which is equivalent to
$\max_{\tilde{\lambda},\kappa}\quad\sum_{i=1}^{K}-y_{i}\log(1+\exp(-\tilde{\lambda}|\theta_{i}-\theta^{*}|^{\kappa}))-(1-y_{i})\log(1+\exp(\tilde{\lambda}|\theta_{i}-\theta_{i}^{*}|^{\kappa}))\,.$
We solve this optimization problem simply by enumerating different values of $\tilde{\lambda}$ and $\kappa$.
Estimation process for $\sigma$. The other group is asked to directly estimate the percentage of total area. For the figure with true parameter values $\theta_{i}^{*}$, the human provides responses by $y_{i}=\theta_{i}^{*}+\epsilon_{i}$. We estimate the variance of the noise by $\hat{\sigma}^{2}=\sum_{i=1}^{N}\frac{(y_{i}-\theta_{i}^{*})^{2}}{N-1}$.
Empirical result. We applied a bootstrap resampling method to the data collected from Amazon MTurk, generating 500 resamples to estimate key parameters, including $\kappa$, $\tilde{\lambda}$, and $\sigma$. The resulting mean estimates are $\kappa=0.328$, $\tilde{\lambda}=0.132$, and $\sigma^{2}=29.665^{2}$. The variances are 0.061, 0.006, and 4.600 for $\kappa$, $\tilde{\lambda}$, and $\sigma$, respectively.
Figure 8 presents the median empirical error of the two-stage framework across varying precision levels, based on 100 repetitions for each value of $\varepsilon$. The results demonstrate that, for any given precision level, the empirical error remains within the targeted threshold. In addition, an upward trend in the empirical error is observed as the precision level increases. The green plot illustrates the sample size required to achieve a specified precision, showing a rapid decline in sample size as precision increases.
For a clearer comparison with the pure SL, Figure 9 displays the error ratio between the two-stage framework and the pure SL. When $\varepsilon=0.3\%$, the median of the estimation error of the two-stage framework is approximately 16% of the pure SL. Increasing the precision to $\varepsilon=4.3\%$, the median of the estimation error of the two-stage framework is approximately 45% of the pure SL. This trend of a decreasing error ratio with tighter precision aligns with Proposition 5.4, highlighting the increasing advantage of the two-stage framework as $\varepsilon$ approaches 0.
Figure 8: Empirical error of the two-stage framework for different values of $\varepsilon$.
Figure 9: Error ratio of the two-stage framework to the pure SL for different values of $\varepsilon$.
In conclusion, we acknowledge limitations in our numerical experiments. Ideally, our two-stage framework could be used to acquire responses sequentially from freelancers, first asking them which number is closer and and then asking them for an estimate of the number of points. Then the two results could be compared. However, implementing this comparison would be challenging because of the adaptive nature of our acquisition process, where each subsequent question depends on previously collected answers. Moreover, acquiring responses from the Amazon MTurk platform is costly. Given the complexity and expense of this procedure, we opted to use simulations to evaluate the performance of our framework.
8 Discussions and Concluding Remarks
This work is motivated by the ever-growing observations of human-AI interaction in LLM. We presented a theoretical two-stage framework for explaining the significance of using human comparisons for model improvement. In the first stage, the noisy-labeled data is fed into the SL procedure to learn the low-dimensional representation; we then ask human evaluators to make pair-wise comparisons in the second stage. To address the challenge of efficient acquisition of valuable information from these human comparisons, we introduced a probabilistic bisection method, which factors in the uncertainty that arises from the accuracy of human comparisons. The newly introduced concept, the LNCA ratio, quantifies the relative scale between label noise and human comparison accuracy. We demonstrate the significant advantages of our proposed two-stage framework over a pure supervised learning approach under certain conditions.
8.1 Implications for Supervised Fine-Tuning
Supervised fine-tuning (SFT) is the process of refining a pre-trained model on a labeled dataset with specific input-output pairs tailored to a target task. Our proposed framework can fit into SFT in the following way. First, the pre-trained model produces a representation, denoted as $\hat{\varphi}(\cdot)$, along with initial parameters $\hat{\boldsymbol{\theta}}_{0}$. For a new task, if $\hat{\varphi}(\cdot)$ remains the same, we proceed directly to the second phase, where human feedback is solicited for model alignment through comparison data. However, if the representation $\hat{\varphi}(\cdot)$ also requires adaptation, we first update the embeddings by training on a combination of newly collected and existing noisy-labeled data, using techniques such as transfer learning (Pan 2020). In this adapted embedding space, human comparisons are then actively acquired in the second stage to enhance model alignment.
8.2 Future Directions
Our work represents an early attempt to organically integrate two sources of data: noisy-labeled data and human comparison data. We position this paper as a prompt for an open discussion. Some potential extensions to our work are worth investigating. First, we extended the probabilistic bisection algorithm from one dimension to multiple dimensions by refining along each dimension. This expansion raises the question of whether a more efficient bisection algorithm exists in multi-dimensional spaces. As discussed in Frazier et al. (2019), the development of a PBA tailored to multi-dimensional problems remains an open problem. Second, although our paper focuses on a prediction problem, the framework has the potential to be extended to action-based learning, where the goal is to select optimal actions for the decision-making problem. Third, exploring the optimal sample selection to maximize the learning algorithm’s efficiency is a compelling area of inquiry. Although we have offered guidelines for practical sample selection, the quest for the most efficient proposed procedure remains open.